### v3

[…]
1 Introduction
Optimal transport (OT) has achieved much success in various areas, such as generative modeling [27, 62, 49, 60], distributional robust optimization [24, 25, 73], non-parametric testing [81, 64, 79, 77, 71], domain adaptation [2, 14, 12, 13, 78], etc. See [63] for comprehensive reviews on these topics. The sample complexity of the Wasserstein distance has been an essential building block for OT in statistical inference. It studies the relationship between a population distribution $\mu$ and its empirical distribution $\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ with $x_{i}\sim\mu$ in terms of the "Wasserstein distance". Unfortunately, the sample size $n$ needs to be exponentially large in data dimension to achieve an accurate enough estimation [22], called the curse of dimensionality issue.
To tackle the challenge of high dimensionality, it is meaningful to combine OT with projection operators in low-dimensional spaces. Researchers first attempted to study the sliced Wasserstein distance [10, 11, 17, 39, 38, 54, 56], which computes the average of the Wasserstein distance between two projected distributions using random one-dimensional projections. Since a single random projection contains little information to distinguish two high-dimensional distributions, computing the sliced Wasserstein distance requires a large number of linear projections. To address this issue, more recent literature considered the Max-Sliced (MS) Wasserstein distance that seeks the optimal projection direction such that the Wasserstein distance between projected distributions is maximized [16, 45, 47, 61, 74]. Later, Wang et al. [76] modified the MS Wasserstein distance by seeking an optimal nonlinear projection belonging to a ball of reproducing kernel Hilbert space (RKHS), which we call the Kernel Max-Sliced (KMS) Wasserstein distance. The motivation is that a nonlinear projector can be more flexible in capturing the differences between two high-dimensional distributions; it is worth noting that KMS Wasserstein reduces to MS Wasserstein when specifying a dot product kernel.
Despite promising applications of the KMS Wasserstein distance, its statistical and computational results have not yet been fully developed. From a statistical perspective, Wang et al. [76] built concentration properties of the empirical KMS Wasserstein distance for distribution that satisfies the Poincar’e inequality projection and the Poincar’e inequality [42], which could be difficult to verify in practice. From a computational perspective, the authors therein developed a gradient-based algorithm to approximately compute the empirical KMS Wasserstein distance. However, there is no theoretical guarantee on the quality of the local optimum solution obtained. In numerical experiments, the quality of the solution obtained is highly sensitive to the initialization.
To address the aforementioned limitations, this paper provides new statistical and computational guarantees for the KMS Wasserstein distance. Our key contributions are summarized as follows.
  • We provide a non-asymptotic estimate on the KMS $p$-Wasserstein distance between two empirical distributions based on $n$ samples, referred to as the finite-sample guarantees. Our result shows that when the samples are drawn from identical populations, the rate of convergence is $n^{-1/(2p)}$, which is dimension-free and optimal in the worst case scenario.
  • We analyze the computation of KMS $2$-Wasserstein distance between two empirical distributions based on $n$ samples. First, we show that computing this distance exactly is NP-hard. Consequently, we are prompted to propose a semidefinite relaxation (SDR) as an approximate heuristic with various guarantees.
    • We develop an efficient first-order method with biased gradient oracles to solve the SDR, the complexity of which for finding a $\delta$-optimal solution is $\widetilde{\mathcal{O}}\left(n^{2}\delta^{-3}\right)$. In comparison, the complexity of the interior point method for solving SDR is $\widetilde{\mathcal{O}}(n^{6.5})$ [5].
    • We derive theoretical guarantees for the optimal solutions from the SDR. We show that there exists an optimal solution from SDR that is at most rank-$k$, where $k\triangleq 1+\lfloor\sqrt{2n+9/4}-3/2\rfloor$, whereas computing the KMS distance exactly requires a rank-$1$ solution. We also provide a corresponding rank reduction algorithm designed to identify such low-rank solutions from the pool of optimal solutions of SDR.
  • We exemplify our theoretical results in non-parametric two-sample testing, human activity detection, and generative modeling. Our numerical results showcase the stable performance and quick computational time of our SDR formulation, as well as the desired sample complexity rate of the empirical KMS Wasserstein distance.
In the following, we compare our work with the most closely related literature, and defer the detailed comparision of KMS Wasserstein distance with other variants of OT divergences in Appendix 8.
Literature Review. The study on the statistical and computational results of MS and KMS Wasserstein distances is popular in the existing literature. From a statistical perspective, existing results on the rate of empirical MS/KMS Wasserstein are either dimension-dependent, suboptimal or require regularity assumptions (e.g., log-concavity, Poincaré inequality, projection Bernstein tail condition) on the population distributions [57, 3, 47, 74], except for the very recent literature [9] that provides a sharp, dimension-free rate for MS Wasserstein with data distributions supported on a compact subspace but without regularity assumptions. From the computational perspective, there are two main approaches to compute such distances. One is to apply gradient-based algorithms to find local optimal solutions or stationary points, see, e.g., [46, 34, 33, 35, 76]. Unfortunately, due to the highly non-convex nature of the optimization problem, the quality of the estimated solution is unstable and highly depends on the choice of initial guess. The other is to consider solving its SDR instead [61], yet theoretical guarantees on the solution from convex relaxation are missing. Inspired by existing reference [4, 18, 59, 44] that studied the rank bound of SDR for various applications, we adopt their proof techniques to provide similar guarantees for computing KMS in Theorem 4.5. Besides, all listed references add entropic regularization to the inner OT problem and solve the regularized version of MS/KMS Wasserstein distances instead, while the gap between the solutions from regularized and original problems could be non-negligible.
2 Background
We first introduce the definition of Wasserstein and KMS Wasserstein distances below.
Definition 2.1 (Wasserstein Distance)
Let $p\in[1,\infty)$. Given a normed space $(\mathcal{V},\|\cdot\|)$, the $p$-Wasserstein distance between two probability measures $\mu,\nu$ on $\mathcal{V}$ is defined as
$W_{p}(\mu,\nu)=\left(\min_{\pi\in\Gamma(\mu,\nu)}\int\|x-y\|^{p}\,\mathrm{d}\pi(x,y)\right)^{1/p},$
where $\Gamma(\mu,\nu)$ denotes the set of all probability measures on $\mathcal{V}\times\mathcal{V}$ with marginal distributions $\mu$ and $\nu$.
Definition 2.2 (RKHS)
Consider a symmetric and positive definite kernel $K:\mathcal{B}\times\mathcal{B}\to\mathbb{R}$, where $\mathcal{B}\subseteq\mathbb{R}^{d}$. Given such a kernel, there exists a unique Hilbert space $\mathcal{H}$, called RKHS, associated with the reproducing kernel $K$. Denote by $K_{x}$ the kernel section at $x\in\mathcal{B}$ defined by $K_{x}(y)=K(x,y),\forall y\in\mathcal{B}$. Any function $f\in\mathcal{H}$ satisfies the reproducing property $f(x)=\langle f,K_{x}\rangle_{\mathcal{H}},\forall x\in\mathcal{B}$. For $x,y\in\mathcal{B}$, it holds that $K(x,y)=\langle K_{x},K_{y}\rangle_{\mathcal{H}}$.
Definition 2.3 (KMS Wasserstein Distance)
Let $p\in[1,\infty)$. Given two distributions $\mu$ and $\nu$, define the $p$-KMS Wasserstein distance as
$\mathcal{KMS}_{p}(\mu,\nu)=\max_{f\in\mathcal{H},~{}\|f\|_{\mathcal{H}}\leq 1}~{}W_{p}(f_{\#}\mu,f_{\#}\nu),$
where $f_{\#}\mu$ and $f_{\#}\nu$ are the pushforward measures of $\mu$ and $\nu$ by $f:\mathcal{B}\to\mathbb{R}$, respectively.
In particular, for dot product kernel $K(x,y)=x^{\mathrm{T}}y$, the RKHS =$\mathcal{H}=\{f:~{}f(x)=x^{\mathrm{T}}\beta,\exists\beta\in\mathbb{R}^{d}\}$. In this case, the KMS Wasserstein distance reduces to the max-sliced Wasserstein distance [16]. A more flexible choice is the Gaussian kernel $K(x,y)=\exp(-\frac{1}{2\sigma^{2}}\|x-y\|^{2}_{2})$, where $\sigma>0$ denotes the temperature hyper-parameter. In this case, the function class $\mathcal{H}$ satisfies the universal property as it is dense in the continuous function class with respect to the $\infty$-type functional norm. It is easy to see the following theorem holds.
Theorem 2.4 (Metric Property of KMS)
Let $\mathcal{H}$ be a universal RKHS. Then $\mathcal{KMS}_{p}(\mu,\nu)=0$ if and only if $\mu=\nu$.
Example 2.5
We present a toy example that highlights the flexibility of the KMS Wasserstein distance. Fig.1(a) displays a scatter plot of the circle dataset, which consists of two groups of samples distributed along inner and outer circles, perturbed by Gaussian noise. Since the data exhibit a nonlinear structure, distinguishing these groups using a linear projection is challenging. As shown in Fig.1(b), the density plot of the projected samples using the MS Wasserstein distance with a linear projector is not sufficiently discriminative. In contrast, Fig. 1(c) demonstrates that the KMS Wasserstein distance is better suited for distinguishing these two groups. Intuitively, the optimal nonlinear projector should take the form $f^{*}(x)\propto\|x\|_{2}^{c}$ for some scalar $c>0$, as illustrated in Fig.1(d). We plot the nonlinear projector by computing the empirical KMS Wasserstein distance, shown in Fig.1(e), which closely resembles the circular landscape depicted in Fig. 1(d). This result demonstrates that the KMS Wasserstein distance provides a data-driven, non-parametric nonlinear projector capable of effectively distinguishing distinct data groups.
Figure 1: Results on a 2-dimensional toy example. (a) Scatter plot of circle dataset; (b) Density plot using MS Wasserstein; (c) Density plot using KMS Wasserstein; (d) Plot of $f(x)=\|x\|_{2},x\in[0,1]^{2}$; (e) Plot of estimated projector using KMS Wasserstein.
Given the RKHS $\mathcal{H}$, let the canonical feature map that embeds data to $\mathcal{H}$ as
$\Phi:~{}\mathcal{B}\to\mathcal{H},\qquad x\mapsto\Phi(x)=K_{x}.$ (1)
Define the functional $u_{f}:~{}\mathcal{H}\to\mathbb{R}$ by $u_{f}(g)=\langle f,g\rangle_{\mathcal{H}}$ for any $g\in\mathcal{H}$, which can be viewed as a linear projector that maps data from the Hilbert space $\mathcal{H}$ to the real line. In light of this, for two probability measures $\mu$ and $\nu$ on $\mathcal{H}$, we define the MS $p$-Wasserstein distance
$\mathcal{MS}_{p}(\mu,\nu)=\sup_{f\in\mathcal{H}:~{}\|f\|_{\mathcal{H}}\leq 1}W_{p}\Big{(}(u_{f})_{\#}\mu,(u_{f})_{\#}\nu\Big{)},$ (2)
where $(u_{f})_{\#}\mu$ denotes the pushforward measure of $\mu$ by the map $u_{f}$, i.e., if $\mu$ is the distribution of a random element $X$ of $\mathcal{H}$, then $(u_{f})_{\#}\mu$ is the distribution of the random variable $u_{f}(X)=\langle f,X\rangle_{\mathcal{H}}$, and $(u_{f})_{\#}\nu$ is defined likewise. In the following, we show that the KMS Wasserstein distance in Definition 2.3 can be reformulated as the MS Wasserstein distance between two distributions on (infinite-dimensional) Hilbert space.
Remark 2.6 (Reformulation of KMS Wasserstein)
By the reproducing property, we can see that $f(x)=\langle f,K_{x}\rangle_{\mathcal{H}}=u_{f}(\Phi(x))$, which implies $f=u_{f}\circ\Phi$. As a consequence,
$\mathcal{KMS}_{p}(\mu,\nu)$ (3)
$=$ $\sup_{f\in\mathcal{H}:~{}\|f\|_{\mathcal{H}}\leq 1}~{}W_{p}\Big{(}(u_{f})_{\#}\big{(}\Phi_{\#}\mu\big{)},(u_{f})_{\#}\big{(}\Phi_{\#}\nu\big{)}\Big{)}$
$=$ $\mathcal{MS}_{p}\Big{(}\Phi_{\#}\mu,\Phi_{\#}\nu\Big{)}.$
In other words, the KMS Wasserstein distance first maps data points into the infinite-dimensional Hilbert space $\mathcal{H}$ through the canonical feature map $\Phi$, and next finds the linear projector to maximally distinguish data from two populations. Compared with the traditional MS Wasserstein distance [16] that performs linear projection in $\mathbb{R}^{d}$, KMS Wasserstein distance is a more flexible notion.
Remark 2.7 (Connections with Kernel PCA)
Given data points $x_{1},\ldots,x_{n}$ on $\mathcal{B}$, denote by $\widehat{\mu}_{n}$ the corresponding empirical distribution. Assume $\frac{1}{n}\sum_{i\in[n]}\Phi(x_{i})=0$, since otherwise one can center those data points as a preprocessing step. Kernel PCA [52] is a popular tool for nonlinear dimensionality reduction. When seeking the first principal nonlinear projection function $f$, Mairal and Vert [51] presents the following reformulation of kernel PCA:
$\underset{f\in\mathcal{H}:~{}\|f\|_{\mathcal{H}}\leq 1}{\arg\max}~{}\text{Var}\Big{(}(u_{f})_{\#}(\Phi_{\#}\widehat{\mu}_{n})\Big{)},$ (4)
where $\text{Var}(\cdot)$ denotes the variance of a given probability measure. In comparison, the KMS Wasserstein distance aims to find the optimal nonlinear projection that distinguishes two populations and replaces the variance objective in (4) with the Wasserstein distance between two projected distributions in (3). Also, kernel PCA is a special case of KMS Wasserstein by taking $p=2,\mu\equiv\widehat{\mu}_{n},\nu\equiv\delta_{0}$ in (3).
[…]
$\Big{\{}\pi\in\mathbb{R}_{+}^{n\times n}:~{}\sum_{i=1}^{n}\pi_{i,j}=\frac{1}{n},\sum_{j=1}^{n}\pi_{i,j}=\frac{1}{n},\forall i,j\in[n]\Big{\}}.$ (5)
Let $\text{Conv}(P)$ denote a convex hull of the set $P$, and $\mathbb{S}_{n}^{+}$ denote the set of positive semidefinite matrices of size $n\times n$. We use $\widetilde{\mathcal{O}}(\cdot)$ as a variant of $\mathcal{O}(\cdot)$ to hide logarithmic factors.
3 Statistical Guarantees
Suppose samples $x^{n}:=\{x_{i}\}_{i\in[n]}$ and $y^{n}:=\{y_{i}\}_{i\in[n]}$ are given and follow distributions $\mu,\nu$, respectively. Denote by $\widehat{\mu}_{n}$ and $\widehat{\nu}_{n}$ the corresponding empirical distributions from samples $x^{n}$ and $y^{n}$. In this section, we provide a finite-sample guarantee on the $p$-KMS Wasserstein distance between $\widehat{\mu}_{n}$ and $\widehat{\nu}_{n}$ with $p\in[1,\infty)$. This guarantee can be helpful for KMS Wasserstein distance-based hypothesis testing that has been studied in [75]: Suppose one aims to build a non-parametric test to distinguish two hypotheses $H_{0}:~{}\mu=\nu$ and $H_{1}:~{}\mu\neq\nu$. Thus, it is crucial to control the high-probability upper bound of $\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})$ under $H_{0}$ as it serves as the critical value to determine whether $H_{0}$ is rejected or not. We first make the following assumption on the kernel. {assumption} There exists some constant $A>0$ such that the kernel $K(\cdot,\cdot)$ satisfies $\sqrt{K(x,x)}\leq A,\forall x\in\mathcal{B}$. Assumption 3 is standard in the literature (see, e.g., [28]), and is quite mild: Gaussian kernel $K(x,y)=\exp(-\|x-y\|_{2}^{2}/\sigma^{2})$ naturally fits into this assumption. For dot product kernel $K(x,y)=x^{\mathrm{T}}y$, if we assume the support $\mathcal{B}$ has a finite diameter, this assumption can also be satisfied. Define the critical value
$\Delta(n,\alpha)=4A\Big{(}C+4\sqrt{\log\frac{2}{\alpha}}\Big{)}^{1/p}\cdot n^{-1/(2p)},$
where $C\geq 1$ is a universal constant. We show the following finite-sample guarantees on KMS $p$-Wassersrein distance.
Theorem 3.1 (Finite-Sample Guarantees)
Fix $p\in[1,\infty)$, level $\alpha\in(0,1)$, and suppose Assumption 3 holds.
  1. (I)
    (One-Sample Guarantee) With probability at least $1-\alpha$, it holds that $\mathcal{KMS}_{p}(\widehat{\mu}_{n},\mu)\leq\frac{1}{2}\Delta(n,\alpha)$.
  2. (II)
    (Two-Sample Guarantee) With probability at least $1-\alpha$, it holds that
    $\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})\leq\mathcal{KMS}_{p}({\mu},{\nu})+\Delta(n,\alpha).$
The dimension-free upper bound $\Delta(n,\alpha)=O(n^{-1(/2p)})$ is optimal in the worst case. Indeed, in the one-dimension case $\mathcal{B}=[0,1]$ and $K(x,y)=xy$, the kernel max-sliced Wasserstein distance $\mathcal{KMS}_{p}$ coincides with the classical Wasserstein distance $W_{p}$. In this case, it is easy to see that if $\mu=(\delta_{0}+\delta_{1})/2$ is supported on the two points 00 and $1$, the expectation of $\mathcal{KMS}(\widehat{\mu}_{n},\widehat{\nu}_{n})$ is of order $n^{-1/(2p)}$ [22]. We also compare this bound with other OT divergences in Appendix 8.
We design a two-sample test $\mathcal{T}_{\mathrm{KMS}}$ such that $H_{0}$ is rejected if $\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})>\Delta(n,\alpha)$. By Theorem 3.1, we have the following performance guarantees of $\mathcal{T}_{\mathrm{KMS}}$.
Corollary 3.2 (Testing Power of $\mathcal{T}_{\mathrm{KMS}}$)
Fix a level $\alpha\in(0,1/2)$, $p\in[1,\infty)$, and suppose Assumption 3 holds. Then the following result holds:
  1. (I)
    (Risk): The type-I risk of $\mathcal{T}_{\mathrm{KMS}}$ is at most $\alpha$;
  2. (II)
    (Power): Under $H_{1}:~{}\mu\neq\nu$, suppose the sample size $n$ is sufficiently large such that $\varrho_{n}:=\mathcal{KMS}_{p}(\mu,\nu)-\Delta(n,\alpha)>0$, the power of $\mathcal{T}_{\mathrm{KMS}}$ is at least $1-c\cdot n^{-1/(2p)}$, where $c$ is a constant depending on $A,C,p,\varrho_{n}$.
Remark 3.3 (Comparision with Maximum Mean Discrepancy (MMD))
MMD has been a popular kernel-based tool to quantify the discrepancy between two probability measures (see, e.g., [28, 23, 37, 66, 67, 6, 53, 48, 69, 72]), which, for any two probability distributions $\mu$ and $\nu$, is defined as
$\text{MMD}(\mu,\nu)=\max_{\begin{subarray}{c}f\in\mathcal{H},\\ ~{}\|f\|_{\mathcal{H}}\leq 1\end{subarray}}~{}\mathbb{E}_{\mu}[f]-\mathbb{E}_{\nu}[f]$ (6)
$=$ $\max_{\begin{subarray}{c}f\in\mathcal{H},\\ ~{}\|f\|_{\mathcal{H}}\leq 1\end{subarray}}~{}\overline{(u_{f})_{\#}\big{(}\Phi_{\#}\mu\big{)}}-\overline{(u_{f})_{\#}\big{(}\Phi_{\#}\nu\big{)}},$
where $\overline{\xi}$ denotes the mean of a given probability measure $\xi$. The empirical (biased) MMD estimator also exhibits dimension-free finite-sample guarantee as in Theorem 3.1: it decays in the order of $\mathcal{O}(n^{-1/2})$, where $n$ is the number of samples. However, the KMS Wasserstein distance is more flexible as it replaces the mean difference objective in (6) by the Wasserstein distance, which naturally incorporates the geometry of the sample space and is suitable for hedging against adversarial data perturbations [24].
4 Computing $2$-KMS Wasserstein Distance
Let $\widehat{\mu}_{n}$ and $\widehat{\nu}_{n}$ be two empirical distributions supported on $n$ points, i.e., $\widehat{\mu}_{n}=\frac{1}{n}\sum_{i}\delta_{x_{i}},\widehat{\nu}_{n}=\frac{1}{n}\sum_{j}\delta_{y_{j}}$, where $\{x_{i}\}_{i},\{y_{j}\}_{j}$ are data points in $\mathbb{R}^{d}$. This section focuses on the computation of $2$-KMS Wasserstein distance between these two distributions. By Definition 2.3 and monotonicity of square root function, it holds that
$\mathcal{KMS}_{2}(\widehat{\mu}_{n},\widehat{\nu}_{n})$ (KMS)
$=$ $\left(\max_{f\in\mathcal{H},~{}\|f\|_{\mathcal{H}}^{2}\leq 1}~{}\left\{\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j\in[n]}\pi_{i,j}|f(x_{i})-f(y_{j})|^{2}\right\}\right)^{1/2},$
where $\Gamma_{n}$ is defined in (5).
Although the outer maximization problem is a functional optimization that contains uncountably many parameters, one can apply the representer theorem (see below) to reformulate Problem (KMS) as a finite-dimensional optimization.
Theorem 4.1 (Theorem 1 in [76])
There exists an optimal solution to (KMS), denoted as $\widehat{f}$, such that for any $z$,
$\widehat{f}(z)=\sum_{i=1}^{n}a_{x,i}K(z,x_{i})-\sum_{i=1}^{n}a_{y,i}K(z,y_{i}),$ (7)
where $a_{x}=(a_{x,i})_{i\in[n]},a_{y}=(a_{y,i})_{i\in[n]}$ are coefficients to be determined.
Define gram matrix $K(x^{n},x^{n})=(K(x_{i},x_{j}))_{i,j\in[n]}$ and other gram matrices $K(x^{n},y^{n}),$ $K(y^{n},x^{n}),K(y^{n},y^{n})$ likewise, then define the concatenation of gram matrics
$G=\begin{pmatrix}K(x^{n},x^{n})&-K(x^{n},y^{n})\\ -K(y^{n},x^{n})&K(y^{n},y^{n})\end{pmatrix}\in\mathbb{R}^{2n\times 2n}.$ (8)
Assume $G$ is positive definite  [^1] In Appendix 10, we provide its sufficient condition. such that it admits the Cholesky decomposition $G^{-1}=UU^{\mathrm{T}}$. By substituting the expression (7) into (KMS) and calculation (see Appendix 11), we obtain the exact reformulation of (KMS):
$\max_{\omega\in\mathbb{R}^{2n}:~{}\|\omega\|_{2}=1}~{}\left\{\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}(M_{i,j}^{\mathrm{T}}\omega)^{2}\right\}.$ (9)
Here, we omit taking the square root of the optimal value of the max-min optimization problem for simplicity of presentation and define the vector $M_{i,j}=U^{\mathrm{T}}M_{i,j}^{\prime}$, where
$M_{i,j}^{\prime}=\begin{pmatrix}(K(x_{i},x_{l})-K(y_{j},x_{l}))_{l\in[n]}\\ (K(y_{j},y_{l})-K(x_{i},y_{l}))_{l\in[n]}\end{pmatrix}\in\mathbb{R}^{2n}.$
Since Problem (9) is a non-convex program, it is natural to question its computational hardness. The following gives an affirmative answer, whose proof is provided in Appendix 12.
Theorem 4.2 (NP-hardness)
Problem (9) is NP-hard for the worst-case instances of $\{M_{i,j}\}_{i,j}$.
The proof idea of Theorem 4.2 is to find an instance of $\{M_{i,j}\}_{i,j}$ that depends on a generic collection of $n$ vectors $\{A_{i}\}_{i}$ such that solving (9) is at least as difficult as solving the fair-PCA problem [65] with rank-$1$ matrices (or fair beamforming problem [68]) and has been proved to be NP-hard [68]. Interestingly, the computational hardness of the MS Wasserstein distance arises from both the data dimension $d$ and the sample size $n$, whereas that of the KMS Wasserstein distance arises from the sample size $n$ only.
To tackle the computational challenge of solving (9), in the subsequent subsections, we present an SDR formula and propose an efficient first-order algorithm to solve it. Next, we analyze the computational complexity of our proposed algorithm and the theoretical guarantees on SDR.
4.1 Semidefinite Relaxation with Efficient Algorithms
We observe the simple reformulation of the objective in (9):
$\sum_{i,j}\pi_{i,j}(M_{i,j}^{\mathrm{T}}\omega)^{2}=\sum_{i,j}\pi_{i,j}\langle M_{i,j}M_{i,j}^{\mathrm{T}},\omega\omega^{\mathrm{T}}\rangle.$
Inspired by this relation, we use the change of variable approach to optimize the rank-$1$ matrix $S=\omega\omega^{\mathrm{T}}$, i.e., it suffices to consider the equivalent reformulation of (9):
$\max_{\begin{subarray}{c}S\in\mathbb{S}_{+}^{2n},\\ \text{Trace}(S)=1,\text{rank}(S)=1\end{subarray}}~{}\left\{F(S)=\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}\langle M_{i,j}M_{i,j}^{\mathrm{T}},S\rangle\right\}.$ (10)
An efficient SDR is to drop the rank-$1$ constraint to consider the semidefinite program (SDP):
$\max_{S\in\mathcal{S}_{2n}}~{}F(S),~{}~{}~{}\text{ where }~{}~{}~{}\mathcal{S}_{2n}=\Big{\{}S\in\mathbb{S}_{+}^{2n}:~{}~{}\text{Trace}(S)=1\Big{\}}.$ (SDR)
Remark 4.3 (Connection with [81])
We highlight that Xie and Xie [81] considered the same SDR heuristic to compute the MS $1$-Wasserstein distance. However, the authors therein apply the interior point method to solve a large-scale SDP, which has expansive complexity $\mathcal{O}(n^{6.5}\text{polylog}(\frac{1}{\delta}))$ (up to $\delta$-accuracy) [5]. In the following, we present a first-order method that exhibits much smaller complexity $\widetilde{\mathcal{O}}\left(n^{2}\delta^{-3}\right)$ in terms of the problem size $n$ (see Theorem 4.4). Besides, theoretical guarantees on the solution from SDR have not been explored in [81], and we are the first literature to provide these results.
The constraint set $\mathcal{S}_{2n}$ is called the Spectrahedron and admits closed-form Bregman projection. Inspired by this, we propose an inexact mirror ascent algorithm to solve (SDR). Its high-level idea is to iteratively construct an inexact gradient estimator and next perform the mirror ascent on iteration points. By properly balancing the trade-off between the bias and cost of querying gradient oracles, this type of algorithm is guaranteed to find a near-optimal solution [31, 32, 29, 30].
We first discuss how to construct supgradient estimators of $F$. By Danskin’s theorem [7],
$\partial F(S)=\text{Conv}\Big{\{}\sum_{i,j}\pi_{i,j}^{*}(S)M_{i,j}M_{i,j}^{\mathrm{T}}:~{}~{}\pi^{*}(S)\in\Pi(S)\Big{\}},$
where $\Pi(S)$ denotes the set of optimal solutions to the following OT problem:
$\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}\langle M_{i,j}M_{i,j}^{\mathrm{T}},S\rangle.$ (11)
The main challenge of constructing a supgradient estimator is to compute an optimal solution $\pi^{*}(S)\in\Gamma(S)$. Since computing an exactly optimal solution is too expensive, we derive its near-optimal estimator, denoted as $\widehat{\pi}$, and practically use the following supgradient estimator:
$v(S)=\sum_{i,j}\widehat{\pi}_{i,j}M_{i,j}M_{i,j}^{\mathrm{T}}.$ (12)
We adopt the stochastic gradient-based algorithm with Katyusha momentum in [83] to compute a $\epsilon$-optimal solution $\widehat{\pi}$ to (11). It achieves the state-of-the-art complexity $\widetilde{\mathcal{O}}\left(n^{2}\epsilon^{-1}\right)$. See the detailed algorithm in Appendix 13. Next, we describe the main algorithm for solving (SDR). Define the (negative) von Neumann entropy $h(S)=\sum_{i\in[2n]}\lambda_{i}(S)\log\lambda_{i}(S)$, where $\{\lambda_{i}(S)\}_{i}$ are the eignevalues of $S$, and define the von Neumann Bregman divergence
$V(S_{1},S_{2})$ $=h(S_{1})-h(S_{2})-\langle S_{1}-S_{2},\nabla h(S_{2})^{\mathrm{T}}\rangle$
$=\text{Trace}(S_{1}\log S_{1}-S_{1}\log S_{2}).$
Iteratively, we update $S_{k+1}$ by performing mirror ascent with constant stepsize $\gamma>0$:
$S_{k+1}=\underset{S\in\mathcal{S}_{2n}}{\arg\max}~{}\gamma\langle v(S_{k}),S\rangle+V(S,S_{k}),$
which admits the following closed-form update:
$\widetilde{S}_{k+1}=\exp\left(\log S_{k}+\gamma v(S_{k})\right),~{}~{}S_{k+1}=\frac{\widetilde{S}_{k+1}}{\text{Trace}(\widetilde{S}_{k+1})}.$ (13)
The general procedure for solving (SDR) is summarized in Algorithm 1.
Algorithm 1 Inexact Mirror Ascent for solving (SDR)
1:  Input: Max iterations $T$, initial guess $S_{1}$, tolerance $\epsilon$, constant stepsize $\gamma$.
2:  for $k=1,\ldots,T-1$ do
3:     Obtain a $\epsilon$-optimal solution (denoted as $\widehat{\pi}$) to (11)
4:     Construct inexact supgradient $v(S_{k})$ by (12)
5:     Perform mirror ascent by (13)
6:  end for
7:  Return $\widehat{S}_{1:T}=\frac{1}{T}\sum_{k=1}^{T}S_{k}$
4.2 Theoretical Analysis
In this subsection, we establish the complexity and performance guarantees for solving (SDR). Since the constraint set $\mathcal{S}_{2n}$ is compact and the objective in (SDR) is continuous, an optimal solution, denoted by $S^{*}$, is guaranteed to exist with a finite optimal value. A feasible solution $\widehat{S}\in\mathcal{S}_{2n}$ is said to be $\delta$-optimal if it satisfies the condition $F(\widehat{S})-F(S^{*})\leq\delta$. Define the constant $C=\max_{i,j}\|M_{i,j}\|_{2}^{2}$.
Theorem 4.4 (Complexity Bound)
Fix the precision $\delta>0$ and specify hyper-parameters
$T=\left\lceil\frac{16C^{2}\log(2n)}{\delta^{2}}\right\rceil,\qquad\epsilon=\frac{\delta}{4},\qquad\gamma=\frac{\log(2n)}{C\sqrt{T}}.$
Then, the complexity of Algorithm 1 for finding $\delta$-optimal solution to (SDR) is $\mathcal{O}\left(C^{3}n^{2}(\log n)^{3/2}\delta^{-3}\right).$
Next, we analyze the quality of the solution to (SDR). Recall the exact reformulation (9) requires that the optimal solution to be rank-$1$ while the tractable relaxation (SDR) does not enforce such a constraint. Therefore, it is of interest to provide theoretical guarantees on the low-rank solution of (SDR), i.e., we aim to find the smallest integer $k\geq 1$ such that there exists an optimal solution to (SDR) that is at most rank-$k$. The integer $k$ is called a rank bound on (SDR), which is characterized in the following theorem.
Figure 2: Diagram of the rank reduction algorithm. Here $\sigma(\cdot)$ denotes the permutation operator on $[n]$, Step 1 can be implemented using the Hungarian algorithm [41], and Step 2-2 finds a direction that lies in the null space of the constraint of Problem (14).
Figure 3: Comparison of SDR-Efficient with the baseline methods SDR-IPM and BCD in terms of computational time and solution quality. The columns, from left to right, correspond to the synthetic Gaussian dataset (100-dimensional), MNIST, and CIFAR-10. The top plots display the computational time, where the $y$-axis is labeled as "unbounded" if the running time exceeds the 1-hour time limit. The bottom plots present the estimated KMS $2$-Wasserstein distance for each method.
Theorem 4.5 (Rank Bound on (SDR))
There exists an optimal solution to (SDR) of rank at most $k\triangleq 1+\left\lfloor\sqrt{2n+\frac{9}{4}}-\frac{3}{2}\right\rfloor$. As a result,
[…]
The trivial rank bound on (SDR) should be $2n$, as the matrix $S$ is of size $2n\times 2n$. Theorem 4.5 provides a novel rank bound that is significantly smaller than the trivial one.
Proof 4.6
Proof Sketch of Theorem 4.5. We first reformulate (SDR) by taking the dual of the inner OT problem:
$\max_{\begin{subarray}{c}S\in\mathcal{S}_{2n}\\ f,g\in\mathbb{R}^{n}\end{subarray}}~{}\left\{\frac{1}{n}\sum_{i=1}^{n}(f_{i}+g_{i}):~{}f_{i}+g_{j}\leq\langle M_{i,j}M_{i,j}^{\mathrm{T}},S\rangle,\forall i,j\right\}.$ (14)
By Birkhoff’s theorem [8] and complementary slackness of OT, there exists an optimal solution of $(f,g)$ such that at most $n$ constraints of (14) are binding, and with such an optimal choice, one can adopt the convex geometry analysis from [43, 44] to derive the desired rank bound for any feasible extreme point of variable $S$. As the set of optimal solutions of (SDR) has a non-empty intersection with the set of feasible extreme points, the desired result holds.$\square$
It is noteworthy that Algorithm 1 only finds a near-optimal solution $\widehat{S}_{1:T}$ of (SDR), which is not guaranteed to satisfy the rank bound in Theorem 4.5. To fill the gap, we develop a rank-reduction algorithm that further converts $\widehat{S}_{1:T}$ to the feasible solution that maintains the desired rank bound. See the general diagram that outlines this algorithm in Fig. 2 and the detailed description in Appendix 17. We also provide its complexity analysis in the following theorem, though in numerical study the complexity is considerably smaller than the theoretical one.
Theorem 4.7
The rank reduction algorithm in Fig. 2 satisfies that $\mathrm{(I)}$ for a $\delta$-optimal solution to (SDR), it outputs another $\delta$-optimal solution with rank at most $k$; $\mathrm{(II)}$ its worst-case complexity is $\mathcal{O}(n^{5})$.
Additionally, by adopting the proof from Luo et al. [50], we show the optimality gap guarantee in Theorem 4.8. Although the approximation ratio seems overly conservative, we find (SDR) has good numerical performance.
Theorem 4.8 (Relaxation Gap of (SDR))
Denote by $\varepsilon=4\cdot(0.33)^{3}$ an universal constant. Then
[…]
Figure 4: Testing power with a controlled type-I error rate of 0.05 across four datasets. Figures from left to right correspond to (a) Gaussian covariance shift, (b) Gaussian mixture distribution shift, (c) MNIST, and (d) CIFAR-10 with distribution abundance changes.
5 Numerical Study
This section presents experiment results for KMS $2$-Wasserstein distance that is solved using SDR with first-order algorithm and rank reduction (denoted as SDR-Efficient). Baseline approaches include the block coordinate descent (BCD) algorithm [76], which finds stationary points of KMS $2$-Wasserstein, and using interior point method (IPM) by off-the-shelf solver cvxpy [19] for solving SDR relaxation (denoted as SDR-IPM). Each instance is allocated a maximum time budget of one hour. All experiments were conducted on a MacBook Pro with an Intel Core i9 2.4GHz and 16GB memory. Unless otherwise stated, error bars are reproduced using $20$ independent trials. Throughout the experiments, we specify the kernel as Gaussian, with the bandwidth being the median of pairwise distances between data points. Other details and extra numerical studies can be found in Appendices 18 and 19.
Computational Time and Solution Quality. We first compare our approach to baseline methods in terms of running time and solution quality. The quality of a given nonlinear projector is assessed by projecting testing data points from two groups and calculating their $2$-Wasserstein distance. The experimental results, shown in the top of Fig. 3, indicate that for small sample size instances, SDR-IPM requires significantly more time than the other two approaches. Additionally, for small sample size instances, the running time of BCD is slightly shorter than that of SDR-Efficient. However, for larger instances, BCD outperforms SDR-Efficient in terms of running time. This observation aligns with our theoretical analysis, which shows that the complexity of SDR is $\tilde{\mathcal{O}}(n^{2}\delta^{-3})$, lower than the complexity of BCD [76], which is $\tilde{\mathcal{O}}(n^{3}\delta^{-3})$. The plots in the bottom of Fig. 3 show the quality of methods SDR-Efficient and BCD. We find the performance of solving SDR outperforms BCD, as indicated by its larger means and smaller variations. One possible explanation is that BCD is designed to find a local optimum solution for the original non-convex problem, making it highly sensitive to the initial guess and potentially less effective in achieving optimal performance.
High-dimensional Hypothesis Testing. Then we validate the performance of KMS $2$-Wasserstein distance for high-dimensional two-sample testing using both synthetic and real datasets. Baseline approaches include the two-sample testing with other statistical divergences, such as (i) Sinkhorn divergence (Sinkhorn Div[27], (ii) maximum mean discrepancy with Gaussian kernel and median bandwidth heuristic (MMD[28], (iii) sliced Wasserstein distance (SW[10], (iv) generalized-sliced Wasserstein distance (GSW[38], (v) max-sliced Wasserstein distance (MS[16], and (vi) optimized mean-embedding test (ME[36]. The baselines KMS, MS, ME partition data into training and testing sets, learn parameters from the training set, and evaluate the testing power on the testing set. Other baselines utilize both sets for evaluation. Synthetic datasets include the high-dimensional Gaussian distributions with covariance shift, or Gaussian mixture distributions. Real datasets include the MNIST [15] and CIFAR-10 [40] with changes in distribution abundance. The type-I error is controlled within $0.05$ for all methods.
We report the testing power for all these approaches and datasets in Fig. 4. For the Gaussian covariance shift scenario the MS method achieves the best performance, which can be explained by the fact that linear mapping is optimal to separate the high-dimensional Gaussian distributions. For the other three scenarios, our approach has the superior performance compared with those baselines.
Table 1: Detection delay of various methods with controlled false alarm rate $\alpha=0.05$. Mean and standard deviation (std) are calculated based on data from 10 different users.
Method KMS Sinkhorn Div MMD SW GSW MS ME
Mean 11.4 16.5 50.6 17.2 12.9 17.8 65.4
Std 5.56 4.4 39.5 8.7 6.4 9.2 25.7
Human Activity Detection. We evaluate the performance of the KMS Wasserstein distance in detecting human activity transitions as quickly as possible using MSRC-12 Kinect gesture dataset [21]. After preprocessing, the dataset consists of 10 users, each with 80 attributes, performing the action throwing/lifting before/after the change-point at time index 600. We employ a sliding window approach [82] with a false alarm rate of $0.01$ to construct the test statistic at each time index, which increases significantly when a change-point is detected. Figure 5 illustrates the test statistics generated by our method compared to baseline approaches. The experimental results, summarized in Table 1, show that our method achieves superior performance.
Figure 5: Top Left: Illustration of sequential data before and after the change-point. Remaining: Testing statistics computed from our and baseline approaches.
Generative Modeling. Finally, we replace the MS Wasserstein distance with our proposed KMS Wasserstein distance for the generative modeling task on the MNIST dataset, following a similar procedure as in [16]. Figure 6 shows the generated samples produced by generative models based on either the sliced Wasserstein distance or the KMS Wasserstein distance. To evaluate the performance of these models, we use the Fréchet Inception Distance (FID) score [heusel2017gans], calculated by extracting image features with a three-layer convolutional neural network. A lower FID score indicates better generative performance. Compared to the baseline, the KMS Wasserstein-based generative model achieves a lower FID score and produces higher-quality images, suggesting that the KMS Wasserstein distance is effective for learning high-dimensional data distributions.
Figure 6: Left: samples generated from sliced Wasserstein-based generative models with FID score 5.37e-2; Right: samples generated from KMS Wasserstein-based generative models with FID score 3.35e-2. Models are trained with feed-forward neural-nets and 30 epoches.
6 Concluding Remarks
In this paper, we presented both statistical and computational guarantees for the KMS Wasserstein distance. From a statistical perspective, we derived finite-sample guarantees under mild conditions on the kernel function, yielding a dimension-free concentration error bound. Moreover, our analysis relaxes the compactness assumption on the sample space typically required by the MS Wasserstein distance. From a computational perspective, we investigated the SDR approach for solving the empirical KMS 2-Wasserstein distance. This method is not only computationally efficient using first-order optimization techniques but also enjoys strong performance guarantees for the obtained solutions. Our numerical experiments validated the theoretical results and demonstrated the superior performance of the KMS Wasserstein distance in high-dimensional hypothesis testing, human activity detection, and generative modeling applications.
Supplementary for “Statistical and Computational Guarantees of Kernel Max-Sliced Wasserstein Distances”
7 Limitations
We list several limitations of this work below.
  1. (I)
    Our theoretical results assume that the sample sizes from sample sets $x^{n}$ and $y^{m}$ satisfy $m=n$ for the simplicity of presentation. If considering the unequal sample size $m\neq n$, the error bound $\Delta(n,\alpha)$ in Theorem 3.1 is replaced by
    $2A\left(C+4\sqrt{\log\frac{1}{\alpha}}\right)^{1/p}\cdot\big{[}n^{-1/(2p)}+m^{-1/(2p)}\big{]};$
    the complexity bound in Theorem 4.4 replaces $n$ with $\max\{n,m\}$; the rank bound in Theorem 4.5 replaces $k=1+\left\lfloor\sqrt{2n+\frac{9}{4}}-\frac{3}{2}\right\rfloor$ with
    $1+\left\lfloor\sqrt{2(n+m-1)+\frac{9}{4}}-\frac{3}{2}\right\rfloor,$
    since for unbalanced OT, at most $n+m-1$ constraints are binding [63, Proposition 3.4].
  2. (II)
    While our statistical guarantees apply to the KMS $p$-Wasserstein distance with general $p\in[1,\infty)$, our computational guarantees are limited to the case of $p=2$. The NP-hardness of computing the KMS $2$-Wasserstein distance suggests that computing the general KMS $p$-Wasserstein distance may also be challenging. This motivates further research into case-specific algorithm design for different values of $p$, e.g., $p=1$.
  3. (III)
    It is desirable to develop exact algorithms or approximation algorithms with better approximation quality for computing KMS $p$-Wasserstein distance.
  4. (IV)
    In this work we only consider classical Gaussian kernel for numerical study. It is desirable to consider deep neural network-based kernel to further boost the power of KMS Wasserstein-based testing.
8 Comparison with Optimal Transport Divergences
Table 2: Time and sample complexity of empirical OT estimators in terms of the number of samples $n$. Here $\widehat{\mu}_{n},\widehat{\nu}_{n}$ represent two empirical distributions based on i.i.d. $n$ sample points in $\mathbb{R}^{d}$. $p$ denotes the order of the metric defined in the cost function of standard Wasserstein distance.
Reference Estimator Name Time Complexity Sample Complexity
[22] $W_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})$ Wasserstein distance $\mathcal{O}(n^{3}\log n)$ $\mathcal{O}(n^{-1/d})$
[26] $\mathcal{S}_{p,\epsilon}(\widehat{\mu}_{n},\widehat{\nu}_{n})$ Sinkhorn Divergence $\mathcal{O}(n^{2})$ per iteration $\mathcal{O}(n^{-1/2}(1+\epsilon^{d/2}))$
[58] $\mathcal{MS}_{k}(\widehat{\mu}_{n},\widehat{\nu}_{n})$ Max Sliced Wasserstein distancewith k-dimensional $\begin{array}[]{c}\text{Max Sliced Wasserstein distance}\\ \text{with $k$-dimensional projector}\end{array}$ $\tilde{\mathcal{O}}(n^{2}d)$ $\mathcal{O}(n^{-1/(\max\{k,2\})})$
[54] $\mathcal{SW}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})$ Sliced Wasserstein distance $\tilde{\mathcal{O}}(nd)$ $\mathcal{O}(n^{-1/2})$
[76] and this work $\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})$ Kernel Max Sliced Wasserstein Distance $\tilde{\mathcal{O}}(n^{2}d^{3})$ $\mathcal{O}(n^{-1/2})$
In Table 2, we summarize the time and sample complexity of various OT divergences. Notably, the sample complexity of the standard Wasserstein distance suffers from the curse of dimensionality. While the Sinkhorn divergence mitigates this issue, its sample complexity depends on $\epsilon^{d/2}$, which can be prohibitively large when the regularization parameter $\epsilon$ is very small. For Max-Sliced (MS) and Sliced Wasserstein distances, their sample complexity is independent of the data dimension. However, they rely on linear projections to analyze data samples, which may not be optimal, particularly when the data exhibits a nonlinear low-dimensional structure. This limitation motivates the study of the KMS Wasserstein distance by Wang et al. [76] and in our work. In Example 2.5 and Section 5, we numerically validate this observation and demonstrate its practical advantages.
9 Proof of Theorem 3.1 and Corollary 3.2
The proof in this part relies on the following technical results.
Theorem 9.1
(Finite-Sample Guarantee on MS $1$-Wasserstein Distance on Hilbert Space, Adopted from [9, Corollary 2.8]) Let $\delta\in(0,1]$, and $\mu$ be a probability measure on a separable Hilbert space $\mathcal{H}$ with $\int_{\mathcal{H}}\|x\|\,\mathrm{d}\mu(x)<\infty$. Let $X_{1},\ldots,X_{n}$ be i.i.d. random elements of $\mathcal{H}$ sampled according to $\mu$, and $\widehat{\mu}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$, then it holds that
$\mathbb{E}\mathcal{MS}_{1}\left(\mu,\widehat{\mu}_{n}\right)\leq C\cdot\left(\int_{\mathcal{H}}\|x\|^{2+2\delta}\,\mathrm{d}\mu(x)\right)^{1/(2+2\delta)}\cdot(\delta n)^{-1/2},$
where $C\geq 1$ is a universal constant.
Theorem 9.2 (Functional Hoefffding Theorem [70, Theorem 3.26])
Let $\mathcal{F}$ be a class of functions, each of the form $h:~{}\mathcal{B}\to\mathbb{R}$, and $X_{1},\ldots,X_{n}$ be samples i.i.d. drawn from $\mu$ on $\mathcal{B}$. For $i\in[n]$, assume there are real numbers $a_{i,h}\leq b_{i,h}$ such that $h(x)\in[a_{i,h},b_{i,h}]$ for any $x\in\mathcal{B},h\in\mathcal{F}\cup\{-\mathcal{F}\}$. Define
$L^{2}=\sup_{h\in\mathcal{F}\cup\{-\mathcal{F}\}}~{}\frac{1}{n}\sum_{i=1}^{n}~{}(b_{i,h}-a_{i,h})^{2}.$
For all $\delta\geq 0$, it holds that
$\mathbb{P}\left\{\sup_{h\in\mathcal{F}}~{}\left|\frac{1}{n}\sum_{i=1}^{n}h(X_{i})\right|\geq\mathbb{E}\left[\sup_{h\in\mathcal{F}}~{}\left|\frac{1}{n}\sum_{1}^{n}h(X_{i})\right|\right]+\delta\right\}\leq\exp\left(-\frac{n\delta^{2}}{4L^{2}}\right).$
We first show the one-sample guarantees for KMS $p$-Wasserstein distance.
Proposition 9.3
Fix $p\in[1,\infty)$, error probability $\alpha\in(0,1)$, and suppose Assumption 3 holds. Let $C\geq 1$ be a universal constant. Then, we have the following results:
  1. (I)
    $\mathbb{E}\mathcal{KMS}_{p}(\mu,\widehat{\mu}_{n})\leq A(2C^{1/p})\cdot n^{-1/(2p)}$
  2. (II)
    With probability at least $1-\alpha$, it holds that
    $\mathcal{KMS}_{p}(\mu,\widehat{\mu}_{n})\leq 2^{1-1/p}A\left(C+4\sqrt{\log\frac{1}{\alpha}}\right)^{1/p}\cdot n^{-1/(2p)}.$
Proof 9.4
Proof of Proposition 9.3. Recall from (3) that
$\mathcal{KMS}_{p}(\mu,\nu)=\mathcal{MS}_{p}\Big{(}\Phi_{\#}\mu,\Phi_{\#}\nu\Big{)}.$
Therefore, it suffices to derive one-sample guarantees for $\mathcal{MS}_{p}\Big{(}\Phi_{\#}\mu,\Phi_{\#}\widehat{\mu}_{n}\Big{)}$.
  1. (I)
    Observe that under Assumption 3, we have
    $A^{2}\geq K(x,x)=\langle K_{x},K_{x}\rangle=\|K_{x}\|_{\mathcal{H}}^{2},$
    and therefore $\|\Phi(x)\|_{\mathcal{H}}=\|K_{x}\|_{\mathcal{H}}\leq A,\forall x\in\mathcal{B}$. In other words, for every probability measure $\mu$ on $\mathcal{B}$, the probability measure $\Phi_{\#}\mu$ is supported on the ball in $\mathcal{H}$ centered at the origin with radius $A$. By Theorem 9.1 with $\delta=1$, we obtain
    $\mathbb{E}\mathcal{KMS}_{1}(\mu,\widehat{\mu}_{n})=\mathbb{E}\mathcal{MS}_{1}\Big{(}\Phi_{\#}\mu,\Phi_{\#}\widehat{\mu}_{n}\Big{)}\leq\frac{AC}{\sqrt{n}}.$
    Since $\Phi_{\#}\mu$ and $\Phi_{\#}\widehat{\mu}_{n}$ are supported on the ball of $\mathcal{H}$ centered at the origin with radius $A$, it holds that
    $\mathcal{MS}_{p}\Big{(}\Phi_{\#}\mu,\Phi_{\#}\widehat{\mu}_{n}\Big{)}\leq\Big{[}\mathcal{MS}_{1}\Big{(}\Phi_{\#}\mu,\Phi_{\#}\widehat{\mu}_{n}\Big{)}\cdot(2A)^{p-1}\Big{]}^{1/p}.$
    In other words,
    $\mathcal{KMS}_{p}(\mu,\widehat{\mu}_{n})\leq\Big{[}\mathcal{KMS}_{1}(\mu,\widehat{\mu}_{n})\cdot(2A)^{p-1}\Big{]}^{1/p}.$ (15)
    It follows that
    $\mathbb{E}\mathcal{KMS}_{p}(\mu,\widehat{\mu}_{n})$ $=\mathbb{E}\mathcal{MS}_{p}\Big{(}\Phi_{\#}\mu,\Phi_{\#}\widehat{\mu}_{n}\Big{)}$
    $\leq\mathbb{E}\Big{[}\mathcal{MS}_{1}\Big{(}\Phi_{\#}\mu,\Phi_{\#}\widehat{\mu}_{n}\Big{)}\cdot(2A)^{p-1}\Big{]}^{1/p}$
    $\leq\left\{\mathbb{E}\Big{[}\mathcal{MS}_{1}\Big{(}\Phi_{\#}\mu,\Phi_{\#}\widehat{\mu}_{n}\Big{)}\cdot(2A)^{p-1}\Big{]}\right\}^{1/p}$
    $\leq\left\{\frac{AC}{\sqrt{n}}\cdot(2A)^{p-1}\right\}^{1/p}=2^{1-1/p}AC^{1/p}\cdot n^{-1/(2p)}.$
  2. (II)
    For the second part, we re-write $\mathcal{KMS}_{1}(\mu,\widehat{\mu}_{n})$ with $\widehat{\mu}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ using the kantorovich dual reformulation of OT:
    $\mathcal{KMS}_{1}(\mu,\widehat{\mu}_{n})=\sup_{\begin{subarray}{c}f\in\mathcal{H}:~{}\|f\|_{\mathcal{H}}\leq 1,\\ g\text{ is $1$-Lipschitz with $g(0)=0$}\end{subarray}}~{}\left|\frac{1}{n}\sum_{i=1}^{n}\Big{(}g(f(x))-\mathbb{E}_{x\sim\mu}[g(f(x))]\Big{)}\right|,$
    where the additional constraint $g(0)=0$ does not impact the optimal value of the OT problem. In other words, one can represent
    $\mathcal{KMS}_{1}(\mu,\widehat{\mu}_{n})=\sup_{h\in\mathfrak{H}}~{}\Big{|}\frac{1}{n}\sum_{i=1}^{n}h(x_{i})\Big{|},$
    where the function class
    =$\mathfrak{H}=\Big{\{}x\mapsto g(f(x))-\mathbb{E}_{x\sim\mu}[g(f(x))]:~{}g\text{ is $1$-Lipschitz with $g(0)=0$},\quad f\in\mathcal{H},\|f\|_{\mathcal{H}}\leq 1\Big{\}}.$
    Consequently, for any $x$,
    $|g(f(x))|=|g(f(x))-g(0)|\leq|f(x)|=|\langle f,K_{x}\rangle_{\mathcal{H}}|\leq\|f\|_{\mathcal{H}}\|K_{x}\|_{\mathcal{H}}\leq A.$
    One can apply Theorem 9.2 with $\mathcal{F}\equiv\mathfrak{H}$, $a_{i,h}\equiv-A-\mathbb{E}_{x\sim\mu}[g(f(x))]$, $b_{i,h}\equiv A-\mathbb{E}_{x\sim\mu}[g(f(x))]$, where $h(x)=g(f(x))-\mathbb{E}_{x\sim\mu}[g(f(x))]$, to obtain
    $\mathbb{P}\Big{\{}\mathcal{KMS}_{1}(\mu,\widehat{\mu}_{n})\geq\mathbb{E}\left[\mathcal{KMS}_{1}(\mu,\widehat{\mu}_{n})\right]+\delta\Big{\}}\leq\exp\left(-\frac{n\delta^{2}}{4(2A)^{2}}\right)=\exp\left(-\frac{n\delta^{2}}{16A^{2}}\right).$
    Or equivalently, the following relation holds with probability at least $1-\alpha$:
    $\mathcal{KMS}_{1}(\mu,\widehat{\mu}_{n})\leq\mathbb{E}\left[\mathcal{KMS}_{1}(\mu,\widehat{\mu}_{n})\right]+4An^{-1/2}\sqrt{\log\frac{1}{\alpha}}\leq An^{-1/2}\left(C+4\sqrt{\log\frac{1}{\alpha}}\right).$
    By the relation (15), we find that with probability at least $1-\alpha$,
    $\mathcal{KMS}_{p}(\mu,\widehat{\mu}_{n})$
    $\leq$ $\Big{[}An^{-1/2}\left(C+4\sqrt{\log\frac{1}{\alpha}}\right)\cdot(2A)^{p-1}\Big{]}^{1/p}=2^{1-1/p}A\left(C+4\sqrt{\log\frac{1}{\alpha}}\right)^{1/p}\cdot n^{-1/(2p)}.$
$\square$
We now complete the proof of Theorem 3.1. By the triangle inequality, with probability at least $1-2\alpha$, it holds that
$\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})$ $\leq\mathcal{KMS}_{p}(\mu,\widehat{\mu}_{n})+\mathcal{KMS}_{p}(\nu,\widehat{\nu}_{n})$
$\leq 2\cdot 2^{1-1/p}A\left(C+4\sqrt{\log\frac{1}{\alpha}}\right)^{1/p}\cdot n^{-1/(2p)}$
$\leq 4A\left(C+4\sqrt{\log\frac{1}{\alpha}}\right)^{1/p}\cdot n^{-1/(2p)}.$
Then, substituting $\alpha$ with $\alpha/2$ gives the desired result.
Proof 9.5
Proof of Corollary 3.2. It remains to show the type-II risk when proving this corollary. In particular,
Type-II Risk $=\mathbb{P}_{H_{1}}\big{\{}\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})<\Delta(n,\alpha)\big{\}}$
$=\mathbb{P}_{H_{1}}\big{\{}\mathcal{KMS}_{p}(\mu,\nu)-\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})\geq\mathcal{KMS}_{p}(\mu,\nu)-\Delta(n,\alpha)\big{\}}$
$\leq\mathbb{P}_{H_{1}}\big{\{}\big{|}\mathcal{KMS}_{p}(\mu,\nu)-\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})\big{|}\geq\mathcal{KMS}_{p}(\mu,\nu)-\Delta(n,\alpha)\big{\}}$
$\leq\frac{\mathbb{E}\big{|}\mathcal{KMS}_{p}(\mu,\nu)-\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})\big{|}}{\mathcal{KMS}_{p}(\mu,\nu)-\Delta(n,\alpha)},$
where the last relation is based on the Markov inequality and the assumption that $\mathcal{KMS}_{p}(\mu,\nu)-\Delta(n,\alpha)>0$. Based on the triangular inequality, we can see that
$\mathbb{E}\big{|}\mathcal{KMS}_{p}(\mu,\nu)-\mathcal{KMS}_{p}(\widehat{\mu}_{n},\widehat{\nu}_{n})\big{|}\leq\mathbb{E}[\mathcal{KMS}_{p}(\mu,\widehat{\mu}_{n})]+\mathbb{E}[\mathcal{KMS}_{p}(\nu,\widehat{\nu}_{n})]\leq 2AC^{1/p}\cdot n^{-1/(2p)}.$
Combining these two upper bounds, we obtain the desired result. $\square$
10 Sufficient Condition for Positive Definiteness of Matrix $G$
To implement our computational algorithm, one needs to assume the gram matrix
$G=[K(x^{n},x^{n}),-K(x^{n},y^{n});-K(y^{n},x^{n}),K(y^{n},y^{n})]$
to be strictly positive definite. By the Lemma on the Schur complement (see, e.g., [5, Lemma 4.2.1]), It can be showed that its necessary and sufficient condition should be
$G^{\prime}=[K(x^{n},x^{n}),K(x^{n},y^{n});K(y^{n},x^{n}),K(y^{n},y^{n})]$
is strictly positive definite. By Wendland [80], this requires our data points $\{x_{1},\ldots,x^{n},y_{1},\ldots,y_{n}\}$ are pairwise distinct and $K(x,y)$ is of the form $K(x,y)=\Phi(x-y),$ with $\Phi(\cdot)$ being continuous, bounded, and its Fourier transform is non-negative and non-vanishing. For instance, Gaussian kernel $K(x,y)=e^{-\|x-y\|_{2}^{2}/\sigma^{2}}$ or Bessel kernel $K(x,y)=(c^{2}+\|x\|_{2}^{2})^{-\beta},x\in\mathbb{R}^{d},\beta>d/2$ satisfies our requirement.
11 Reformulation for $2$-KMS Wasserstein Distance in (KMS)
In this section, we derive the reformulation for computing $2$-KMS Wasserstein distance:
$\max_{f\in\mathcal{H},~{}\|f\|_{\mathcal{H}}^{2}\leq 1}~{}\left\{\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j\in[n]}\pi_{i,j}|f(x_{i})-f(y_{j})|^{2}\right\}.$ (16)
Based on the expression of $f$ in (7), we reformulate the problem above as
$\max_{a_{x},a_{y}\in\mathbb{R}^{n}}~{}\left\{\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j\in[n]}\pi_{i,j}\left|\sum_{l\in[n]}a_{x,l}K(x_{i},x_{l})-\sum_{l\in[n]}a_{y,l}K(y_{j},y_{l})\right|^{2}\right\},$ (17a)
subject to the constraint
$\left\|\sum_{i=1}^{n}a_{x,i}K(\cdot,x_{i})-\sum_{i=1}^{n}a_{y,i}K(\cdot,y_{i})\right\|_{\mathcal{H}}^{2}$ (17b)
$=$ $\langle{\sum_{i=1}^{n}a_{x,i}K(\cdot,x_{i})-\sum_{i=1}^{n}a_{y,i}K(\cdot,y_{i})}{\sum_{i=1}^{n}a_{x,i}K(\cdot,x_{i})-\sum_{i=1}^{n}a_{y,i}K(\cdot,y_{i})}\rangle$
$=$ $\sum_{i,j\in[n]}~{}a_{x,i}a_{x,j}\langle K(\cdot,x_{i}),K(\cdot,x_{j})\rangle+\sum_{i,j\in[n]}~{}a_{y,i}a_{y,j}\langle K(\cdot,y_{i}),K(\cdot,y_{j})\rangle-2\sum_{i,j\in[n]}~{}a_{x,i}a_{y,j}\langle K(\cdot,x_{i}),K(\cdot,y_{j})\rangle\leq 1.$
One can re-write (17) in compact matrix form. If we define
$s$ $=[a_{x};a_{y}],$
$M_{i,j}^{\prime}$ $=[(K(x_{i},x_{l})-K(y_{i},x_{l}))_{l\in[n]};(K(y_{j},y_{l})-K(x_{i},y_{l}))_{l\in[n]}],$
$G$ $=[K(x^{n},x^{n}),-K(x^{n},y^{n});-K(y^{n},x^{n}),K(y^{n},y^{n})]\in\mathbb{R}^{2n\times 2n},$
Problem (17) can be reformualted as
$\max_{s\in\mathbb{R}^{2n}}~{}\left\{\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j\in[n]}\pi_{i,j}\left|s^{\mathrm{T}}M_{i,j}^{\prime}\right|^{2}:~{}~{}~{}~{}s^{\mathrm{T}}Gs\leq 1\right\}.$ (18)
Take Cholesky decomposition $G^{-1}=UU^{\mathrm{T}}$ and use the change of variable approach to take $\omega=U^{-1}s$, Problem (18) can be further reformulated as
$\max_{\omega\in\mathbb{R}^{2n}}~{}\left\{\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j\in[n]}\pi_{i,j}\big{(}\langle\omega,U^{\mathrm{T}}M_{i,j}^{\prime}\rangle\big{)}^{2}:~{}~{}~{}~{}\omega^{\mathrm{T}}\omega\leq 1\right\}.$ (19)
After defining $M_{i,j}=U^{\mathrm{T}}M_{i,j}^{\prime}$ and observing that the inequality constraint $\omega^{\mathrm{T}}\omega\leq 1$ will become tight, we obtain the desired reformulation as in (9).
12 Proof of Theorem 4.2
The general procedure of NP-hardness proof is illustrated in the following diagram: Problem (9) contains the (Fair PCA with rank-$1$ data) as a special case, whereas this special problem further contains (Partition) (which is known to be NP-complete) as a special case. After building these two reductions, we finish the proof of Theorem 4.2.
Figure 7: Proof outline of Theorem 4.2
Claim 1. Problem (9) contains Problem (Fair PCA with rank-1 data).
Proof 12.1
Proof of Claim 1. Given vectors $A_{1},\ldots,A_{n}$, we specify
$M_{1,:}\triangleq\{M_{1,1},M_{1,2},\ldots,M_{1,n}\}=\{A_{1},\ldots,A_{n}\},$
and $M_{i,:}\triangleq\{M_{i,1},M_{i,2},\ldots,M_{i,n}\},i=2,\ldots,n$ is specified by circularly shifting the elements in $M_{1,:}$ by $i-1$ positions. For instance, $M_{2,:}=\{A_{n},A_{1},\ldots,A_{n-1}\}$. For the inner OT problem in (9), it suffices to consider deterministic optimal transport $\pi$, i.e.,
$\pi_{i,j}=\left\{\begin{aligned} 1/n,&\quad\text{if $j=\sigma(i)$},\\ 0,&\quad\text{otherwise}\end{aligned}\right.$
for some bijection mapping $\sigma:~{}[n]\to[n]$. The cost matrix for the inner OT is actually a circulant matrix:
$\Big{(}(M_{i,j}^{\mathrm{T}}\omega)^{2}\Big{)}_{i,j}=\begin{pmatrix}(A_{1}\omega)^{2}&(A_{2}\omega)^{2}&\cdots&(A_{n}\omega)^{2}\\ (A_{n}\omega)^{2}&(A_{1}\omega)^{2}&\cdots&(A_{n-1}\omega)^{2}\\ \vdots&\vdots&\ddots&\vdots\\ (A_{2}\omega)^{2}&(A_{3}\omega)^{2}&\cdots&(A_{1}\omega)^{2}\end{pmatrix}.$
When considering the feasible circularly shifting bijection mapping (e.g., $\sigma(i)=(i+j)\mod n,\forall i\in[n]$ for $j=0,1,\ldots,n-1$), we obtain the upper bound on the optimal value of the inner OT problem in (9):
$\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}(M_{i,j}^{\mathrm{T}}\omega)^{2}\leq\min_{i\in[n]}~{}(A_{i}^{\mathrm{T}}\omega)^{2}=\min_{i\in[n]}~{}\omega^{\mathrm{T}}A_{i}A_{i}^{\mathrm{T}}\omega.$
On the other hand, for any bijection mapping $\sigma$, the objective of the inner OT problem in (9) can be written as a convex combination of $(A_{1}^{\mathrm{T}}\omega)^{2},\ldots,(A_{n}^{\mathrm{T}}\omega)^{2}$, and thus,
$\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}(M_{i,j}^{\mathrm{T}}\omega)^{2}\geq\min_{\alpha\in\mathbb{R}_{n}^{+},\sum_{i}\alpha_{i}=1}~{}\Big{\{}\sum_{i}\alpha_{i}(A_{i}^{\mathrm{T}}\omega)^{2}\Big{\}}\geq\min_{i\in[n]}~{}(A_{i}^{\mathrm{T}}\omega)^{2}.$
Since the upper and lower bounds match with each other, we obtain
$\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}(M_{i,j}^{\mathrm{T}}\omega)^{2}=\min_{i\in[n]}~{}\omega^{\mathrm{T}}A_{i}A_{i}^{\mathrm{T}}\omega,$
and consequently,
$\max_{\omega:~{}\|\omega\|_{2}=1}~{}\left\{\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}(M_{i,j}^{\mathrm{T}}\omega)^{2}\right\}=\max_{\omega:~{}\|\omega\|_{2}=1}~{}\left\{\min_{i\in[n]}~{}\omega^{\mathrm{T}}A_{i}A_{i}^{\mathrm{T}}\omega\right\},$
which justifies Problem (9) contains Problem (Fair PCA with rank-1 data). $\square$
Claim 2. Problem (Fair PCA with rank-1 data) contains Problem (Partition).
It is noteworthy that Claim 2 has previously been proved by [68]. For the sake of completeness, we provide the proof here.
Proof 12.2
Proof of Claim 2. Consider the norm minimization problem
$P=\min_{\omega}~{}\Big{\{}\|\omega\|_{2}^{2}:~{}\min_{i\in[n]}(\omega^{\mathrm{T}}A_{i})^{2}\geq 1\Big{\}}.$ (20)
and the scaled problem
$\max_{\omega}~{}\Big{\{}\min_{i\in[n]}~{}(\omega^{\mathrm{T}}A_{i})^{2}:~{}\|\omega\|_{2}^{2}=P\Big{\}}.$ (21)
We can show that Problem (Fair PCA with rank-1 data) is equivalent to (21), whereas (21) is equivalent to (20). Indeed,
  • For the first argument, for any optimal solution from Problem (Fair PCA with rank-1 data), denoted as $\omega^{*}$, one can do the scaling to consider $\tilde{\omega}^{*}=\sqrt{P}\omega$, which is also optimal to (21), and vise versa.
  • For the second argument, let $\omega_{1},\omega_{2}$ be optimal solutions from (20), (21), respectively. Since $P$ is the optimal value of (20), one can check that $\omega_{1}$ is a feasible solution to (21). Since $\min_{i\in[n]}~{}(\omega_{1}^{\mathrm{T}}A_{i})^{2}\geq 1$, by the optimality of $\omega_{2}$, it holds that $\min_{i\in[n]}~{}(\omega_{2}^{\mathrm{T}}A_{i})^{2}\geq 1$, i.e., $\omega_{2}$ is a feasible solution to (20). Since $\|\omega_{2}\|_{2}^{2}=P$, $\omega_{2}$ is an optimal solution to (20). Reversely, one can show $\omega_{1}$ is an optimal solution to (21): suppose on the contrary that there exists $\bar{\omega}_{1}$ such that $\|\bar{\omega}_{1}\|_{2}^{2}=P$ and $\min_{i\in[n]}~{}(\bar{\omega}_{1}^{\mathrm{T}}A_{i})^{2}>\min_{i\in[n]}~{}(\omega_{1}^{\mathrm{T}}A_{i})^{2}\geq 1$, then one can do a scaling of $\bar{\omega}_{1}$ such that $\min_{i\in[n]}~{}(\bar{\omega}_{1}^{\mathrm{T}}A_{i})^{2}=1$ whereas $\|\bar{\omega}_{1}\|_{2}^{2}>P$, which contradicts to the optimality of $P$. Combining both directions, we obtain the equivalence argument.
Thus, it suffices to show (20) contains Problem (Partition). Define $a=(a_{i})_{i\in[n]},Q=I_{n}+aa^{\mathrm{T}}$, and assume $Q$ admits Cholesky factorization $Q=S^{\mathrm{T}}S$. Then we create the vector $A_{i}=S^{-\top}e_{i}$, where $e_{i}$ is the $i$-th unit vector of length $n$. Then, it holds that
[…]
$=$ $\min_{\omega}~{}\Big{\{}\|\omega\|_{2}^{2}:~{}\min_{i\in[n]}((S^{-1}\omega)^{\mathrm{T}}e_{i})^{2}\geq 1\Big{\}}$
$=$ $\min_{\omega}~{}\Big{\{}\|Sx\|_{2}^{2}:~{}\min_{i\in[n]}(x^{\mathrm{T}}e_{i})^{2}\geq 1\Big{\}}$
$=$ $\min_{\omega}~{}\Big{\{}x^{\mathrm{T}}Qx:~{}x_{i}^{2}\geq 1\Big{\}}$
$=$ $\min_{\omega}~{}\Big{\{}\sum_{i=1}^{n}x_{i}^{2}+\left(\sum_{i=1}^{n}a_{i}x_{i}\right)^{2}:~{}x_{i}^{2}\geq 1\Big{\}}\qquad\qquad\qquad(*)$
where the second equality is by the change of variable $x=S^{-1}\omega$, the third equality is by the definitions of $S$ and $e_{i}$, and the last equality is by the definition of $Q$. The solution to Problem (Partition) exists if and only if the optimal value to Problem (*) equals $n$. Thus, we finish the proof of Claim 2. $\square$
13 Algorithm that Finds Near-optimal Solution to Optimal Transport
In this section, we present the algorithm that returns $\epsilon$-optimal solution to the following OT problem:
$\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}c_{i,j},$ (22)
where $\{c_{i,j}\}_{i,j}$ is the given cost matrix. Define $\|C\|_{\infty}=\max_{i,j}c_{i,j}$. In other words, we find $\widehat{\pi}\in\Gamma_{n}$ such that
[…]
Entropy-Regularized OT.
The key to the designed algorithm is to consider the entropy regularized OT problem
$\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}c_{i,j}+\eta\sum_{i,j}\pi_{i,j}\log(\pi_{i,j}),$
whose dual problem is
$\min_{v\in\mathbb{R}^{n}}~{}\left\{G(v)=\frac{1}{n}\sum_{i=1}^{n}h_{i}(v)\right\},$ (23)
where
$h_{i}(v)=\eta\log\sum_{j}\exp\left(\frac{v_{j}-c_{i,j}-\eta}{\eta}\right)-\frac{1}{n}\sum_{j}v_{j}+\eta(1+\log n).$
Given the dual variable $v\in\mathbb{R}^{n}$, one can recover the primal variable $\pi$ using
$\pi(v)=\frac{\frac{1}{n}\exp\left(\frac{v_{j}-c_{i,j}-\eta}{\eta}\right)}{\sum_{j^{\prime}\in[n]}\exp\left(\frac{v_{j^{\prime}}-c_{i,j^{\prime}}-\eta}{\eta}\right)}$
Algorithm 2 essentailly optimizes the dual formulation (23) with light computational speed.
Algorithm 2 Stochastic Gradient-based Algorithm with Katyusha Momentum for solving OT [83]
1:  Input: Accuracy $\epsilon>0$, $\eta=\frac{\epsilon}{8\log n}$, $\epsilon^{\prime}=\frac{\epsilon}{6\max_{i,j}c_{i,j}}$, maximum outer iteration $T_{\text{out}}$, and maximum inner iteration $T$.
2:  Take $(y_{0},z_{0},\tilde{\lambda}_{0},\lambda_{0},C_{0},D_{0})=(0,0,0,0,0,0)$
3:  for $t=0,\ldots,T_{\text{out}}-1$ do
4:     $\tau_{1,t}=\frac{2}{t+4},\gamma_{t}=\frac{\eta}{9\tau_{1,t}}$
5:     $u_{t}=\nabla\phi(\widetilde{\lambda}_{t})$
6:     for $j=0,\ldots,T-1$ do
7:        $k=j+tT$
8:        $\lambda_{k+1}=\tau_{1,t}z_{k}+\frac{1}{2}\widetilde{\lambda}_{t}+(\frac{1}{2}-\tau_{1,t})y_{k}$
9:        Sample $i$ uniformly from $[n]$, and construct
$H_{k+1}=u_{t}+\Big{(}\nabla h_{i}(\lambda_{k+1})-\nabla h_{i}(\widetilde{\lambda}_{t})\Big{)}$
10:        Update $z_{k+1}=z_{k}-\gamma_{t}\cdot H_{k+1}/2$ and $y_{k+1}=\lambda_{k+1}-\eta H_{k+1}/9$
11:     end for
12:     Update $\widetilde{\lambda}_{t+1}=\frac{1}{T}\sum_{j=1}^{T}y_{tT+j}$
13:     Sample $\widehat{\lambda}_{t}$ uniformly from $\{\lambda_{tT+1},\ldots,\lambda_{tT+T}\}$ and take $D_{t}=D_{t}+\texttt{vec}(\pi(\widehat{\lambda}_{t}))/\tau_{1,t}$
14:     $C_{t}=C_{t}+1/\tau_{1,t}$
15:     $\pi_{t+1}=D_{t}/C_{t}$
16:  end for
17:  Query Algorithm 3 to Round $\widetilde{\pi}:=\pi_{T_{\text{out}}}$ to $\widehat{\pi}$ such that $\widehat{\pi}\textbf{1}_{n}=\frac{1}{n}\textbf{1}_{n}$ and $\widehat{\pi}^{\mathrm{T}}\textbf{1}_{n}=\frac{1}{n}\textbf{1}_{n}$
18:  Return $\widehat{\pi}$
Algorithm 3 Round to $\Gamma_{n}$ ([1, Algorithm 2])
1:  Input: $\pi\in\mathbb{R}_{+}^{n\times n}$
2:  $X=\text{diag}(x_{1},\ldots,x_{n})$, with $x_{i}=\min\{1,\frac{1}{nr_{i}(\pi)}\}$. Here $r_{i}(\pi)$ denotes the $i$-th row sum of $\pi$.
3:  $\pi^{\prime}=X\pi$.
4:  $Y=\text{diag}(y_{1},\ldots,y_{n})$, with $y_{j}=\min\{1,\frac{1}{nc_{i}(\pi^{\prime})}\}$. Here $c_{j}(\pi^{\prime})$ denotes the $j$-th column sum of $\pi^{\prime}$.
5:  $\pi^{\prime\prime}=\pi^{\prime}Y$.
6:  $\mathbf{e}_{r}=\frac{1}{n}\textbf{1}_{n}-r(\pi^{\prime\prime}),\mathbf{e}_{c}=\frac{1}{n}\textbf{1}_{n}-c(\pi^{\prime\prime})$, where
$r(\pi^{\prime\prime})=(r_{i}(\pi^{\prime\prime}))_{i\in[n]},c(\pi^{\prime\prime})=(c_{j}(\pi^{\prime\prime}))_{j\in[n]}.$
7:  Return $\pi^{\prime\prime}+\mathbf{e}_{r}\mathbf{e}_{c}^{\mathrm{T}}/\|\mathbf{e}_{r}\|_{1}$.
Theorem 13.1 ([83, Theorem 3])
Suppose we specify $T_{\text{out}}=\mathcal{O}(\frac{\|C\|_{\infty}\sqrt{\ln n}}{\epsilon}),T=n,$ the number of total iterations (including outer and inner iterations) of Algorithm 2 is $\mathcal{O}(\frac{n\|C\|_{\infty}\sqrt{\ln n}}{\epsilon})$ with per-iteration cost $\mathcal{O}(n)$. Therefore, the number of arithmetic operations of Algorithm 2 for finding $\epsilon$-optimal solution is $\mathcal{O}(\frac{n^{2}\|C\|_{\infty}\sqrt{\ln n}}{\epsilon})$
14 Proof of Theorem 4.4
To analyze the complexity of Algorithm 1, we first derive the bias and computational cost of the supgradient estimator $v(S)$ in (12).
Lemma 14.1 (Bias and Computational Cost)
[…]
Next, we analyze the error of the inexact mirror ascent framework in Algorithm 1.
Lemma 14.2 (Error Analysis of Algorithm 1)
When taking the stepsize $\gamma=\frac{\log(2n)}{C\sqrt{T}}$, the output $\widehat{S}_{1:T}$ from Algorithm 1 satisfies
$0\leq F(S^{*})-F(\widehat{S}_{1:T})\leq 2\epsilon+2C\sqrt{\frac{\log(2n)}{T}}.$
Combining Lemmas 14.1 and 14.2, we obtain the complexity for solving (SDR).
Proof 14.3
Proof of Lemma 14.1. For the first part, it is noteworthy that $v(S)$ is associated with the objective
$\widehat{F}(S)=\sum_{i,j}\widehat{\pi}_{i,j}\langle M_{i,j}^{\mathrm{T}}M_{i,j},S\rangle,$
where $\widehat{\pi}_{i,j}$ is the $\epsilon$-optimal solution to
$F(S)=\min\limits_{\pi\in\Gamma_{n}}\sum_{i,j}\pi_{i,j}\langle M_{i,j}^{\mathrm{T}}M_{i,j},S\rangle.$
By definition, it holds that
$0\leq\widehat{F}(S)-F(S)\leq\epsilon.$
The second part follows from Theorem 13.1. $\square$
The proof of Lemma 14.2 replies on the following technical result.
Lemma 14.4 ([55])
Let $\{S_{k}\}_{k=1}^{T}$ be the updating trajectory of mirror ascent aiming to solve the maximization of $G(S)$ with $S\in\mathcal{S}_{2n}$, i.e.,
$S_{k+1}=\underset{S\in\mathcal{S}_{2n}}{\arg\max}~{}\gamma\langle v(S_{k}),S\rangle+V(S,S_{k}),\quad k=1,\ldots,T-1.$
Here $v(S)$ is a supgradient of $G(S)$, and we assume there exists $M_{*}>0$ such that
$\left\|v(S)\right\|_{\text{Tr}}:=\text{Trace}(v(S))\leq M_{*},\qquad\forall S\in\mathcal{S}_{2n}.$
Let $\widehat{S}_{1:T}=\frac{1}{T}\sum_{k=1}^{T}S_{k}$, and $S^{*}$ be a maximizer of $G(S)$. Define the diameter
$D_{\mathcal{S}_{2n}}^{2}=\max_{S\in\mathcal{S}_{2n}}h(S)-\min_{S\in\mathcal{S}_{2n}}h(S)=\log(2n).$
For constant step size
$\gamma=\frac{D_{\mathcal{S}_{2n}}^{2}}{M_{*}\sqrt{T}}=\frac{\log(2n)}{M_{*}\sqrt{T}},$
it holds that
$0\leq G(S^{*})-G(\widehat{S}_{1:T})\leq M_{*}\sqrt{\frac{4\log(2n)}{T}}.$
Proof 14.5
Proof of Lemma 14.2. Let $S^{*}$ and $\widehat{S}^{*}$ be maximizers of the objective $F(\cdot)$ and $\widehat{F}(\cdot)$, then we have the following error decomposition:
$F(S^{*})-F(\widehat{S}_{1:T})$
$=$ $\big{[}F(S^{*})-\widehat{F}(S^{*})\big{]}+\big{[}\widehat{F}(S^{*})-\widehat{F}(\widehat{S}^{*})\big{]}+\big{[}\widehat{F}(\widehat{S}^{*})-\widehat{F}(\widehat{S}_{1:T})\big{]}+\big{[}\widehat{F}(\widehat{S}_{1:T})-F(\widehat{S}_{1:T})\big{]}$
$\leq$ $2\epsilon+\big{[}\widehat{F}({S}^{*})-\widehat{F}(\widehat{S}^{*})\big{]}+\big{[}\widehat{F}(\widehat{S}^{*})-\widehat{F}(\widehat{S}_{1:T})\big{]}$
$\leq$ $2\epsilon+\big{[}\widehat{F}(\widehat{S}^{*})-\widehat{F}(\widehat{S}_{1:T})\big{]},$
where the first inequality is because $\|F-\widehat{F}\|_{\infty}\leq\epsilon$ and
$|\big{[}F(S^{*})-\widehat{F}(S^{*})\big{]}|\leq\epsilon,|\big{[}\widehat{F}(\widehat{S}_{1:T})-F(\widehat{S}_{1:T})\big{]}|\leq\epsilon;$
and the second inequality is because $\widehat{F}(\widehat{S}^{*})-\widehat{F}(\widehat{S}_{1:T})\leq 0$. It remains to bound $\big{[}\widehat{F}(\widehat{S}^{*})-\widehat{F}(\widehat{S}_{1:T})\big{]}$. It is worth noting that
$\|v(S)\|_{\text{Tr}}=\sum_{i,j}\pi_{i,j}\|M_{i,j}M_{i,j}^{\mathrm{T}}\|_{\text{Tr}}\leq\sum_{i,j}\pi_{i,j}\cdot C=C.$
Therefore, the proof can be finished by querying Lemma 14.4 with $M_{*}=C$ and stepsize $\gamma=\frac{\log(2n)}{C\sqrt{T}}$. $\square$
Proof 14.6
Proof of Theorem 4.4. The proof can be finished by taking hyper-parameters such that
$2\epsilon\leq\frac{\delta}{2},\qquad 2C\sqrt{\frac{\log(2n)}{T}}\leq\frac{\delta}{2}.$
In other words, we take $\epsilon=\frac{\delta}{4}$ and $T=\lceil\frac{16C^{2}\log(2n)}{\delta^{2}}\rceil$. We follow the proof of Lemma 14.2 to take stepsize $\gamma=\frac{\log(2n)}{C\sqrt{T}}$$\square$
15 Proof of Theorems 4.5 and 4.8
We rely on the following two technical results when proving Theorem 4.5.
Theorem 15.1 (Birkhoff-von Neumann Theorem [8])
Consider the discrete OT problem
$\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}c_{i,j},.$ (24)
There exists an optimal solution $\pi$ that has exactly one entry of $1/n$ in each row and each column with all other entries 00.
Theorem 15.2 (Rank Bound, Adopted from [44, Theorem 2] and [43, Lemma 1])
Consider the domain set
$\mathcal{D}=\Big{\{}S\in\mathbb{S}_{m}^{+}:~{}\text{Trace}(S)=1\Big{\}}$
and the intersection of $N$ linear inequalities:
=$\mathcal{E}=\Big{\{}S\in\mathbb{R}^{m\times m}:~{}\langle S,A_{i}\rangle\geq b_{i},i\in[N]\Big{\}}.$
Then, any feasible extreme point in $\mathcal{D}\cap\mathcal{E}$ has a rank at most $1+\lceil\sqrt{2N+9/4}-3/2\rceil$. Such a rank bound can be strengthened by replacing $N$ by the number of binding constraints in $\mathcal{E}$.
Proof 15.3
Proof of Theorem 4.5. By taking the dual of inner OT problem, we find (SDR) can be reformulated as
$\max_{\begin{subarray}{c}S\in\mathcal{S}_{2n}\\ f,g\in\mathbb{R}^{n}\end{subarray}}~{}\left\{\frac{1}{n}\sum_{i=1}^{n}(f_{i}+g_{i}):\quad f_{i}+g_{j}\leq\langle M_{i,j}M_{i,j}^{\mathrm{T}},S\rangle,\quad\forall i,j\in[n]\right\}.$ (25)
Let $S^{*}$ be the optimal solution of variable $S$ to the optimization problem above. Then for fixed $S^{*}$, according to Theorem 15.1 and complementary slackness of OT, there exists optimal solutions $(f^{*},g^{*})$ such that only $n$ constraints out of $n^{2}$ constraints in (25) are binding. Moreover, an optimal solution to (SDR) can be obtained by finding a feasible solution to the following intersection of constraints:
Find $\text{Find }S\in\mathcal{S}_{2n}\bigcap\mathcal{E}\triangleq\Big{\{}S:~{}f_{i}^{*}+g_{j}^{*}\leq\langle M_{i,j}M_{i,j}^{\mathrm{T}},S\rangle,~{}~{}i,j\in[n]\Big{\}}.$
By Theorem 15.2, any feasible extreme point from $\mathcal{S}_{2n}\cap\mathcal{E}$ has rank at most $1+\left\lfloor\sqrt{2n+\frac{9}{4}}-\frac{3}{2}\right\rfloor$. Thus, we pick such a feasible extreme point to satisfy the requirement of Theorem 4.5. $\square$
Proof 15.4
Proof of Theorem 4.8. Recall that
(KMS)$=\max_{\begin{subarray}{c}S\succeq 0,\text{Trace}(S)=1,\text{rank}(S)=1,\\ f,g\in\mathbb{R}^{n}\end{subarray}}~{}\left\{\frac{1}{n}\sum_{i=1}^{n}(f_{i}+g_{i}):\quad f_{i}+g_{j}\leq\langle M_{i,j}M_{i,j}^{\mathrm{T}},S\rangle,\quad\forall i,j\in[n]\right\}$
and
(SDR)$=\max_{\begin{subarray}{c}S\succeq 0,\text{Trace}(S)=1\\ f,g\in\mathbb{R}^{n}\end{subarray}}~{}\left\{\frac{1}{n}\sum_{i=1}^{n}(f_{i}+g_{i}):\quad f_{i}+g_{j}\leq\langle M_{i,j}M_{i,j}^{\mathrm{T}},S\rangle,\quad\forall i,j\in[n]\right\}.$
[…]
$\zeta\sim\mathcal{N}(0,\widehat{S}),\qquad\xi=\frac{\zeta}{\|\zeta\|_{2}},\qquad\widetilde{S}=\xi\xi^{\mathrm{T}}.$
It can be seen that $\widetilde{S}\succeq 0,\text{Trace}(\widetilde{S})=1,\text{rank}(\widetilde{S})=1$. Then, for any $\varepsilon\in(0,1]$ and $\mu>0$, it holds that
$\text{Pr}\Big{\{}\langle M_{i,j}M_{i,j}^{\mathrm{T}},\widetilde{S}\rangle\geq\varepsilon\cdot\big{(}\widehat{f}_{i}+\widehat{g}_{j}\big{)},\quad\forall i,j\in[n]\Big{\}}$
$\geq$ $\text{Pr}\Big{\{}\langle M_{i,j}M_{i,j}^{\mathrm{T}},\widetilde{S}\rangle\geq\varepsilon\cdot\langle M_{i,j}M_{i,j}^{\mathrm{T}},\widehat{S}\rangle,\quad\forall i,j\in[n]\Big{\}}$
$=$ $\text{Pr}\Big{\{}\langle M_{i,j}M_{i,j}^{\mathrm{T}},\widetilde{S}\rangle\geq\varepsilon\cdot\mathbb{E}[\langle M_{i,j}M_{i,j}^{\mathrm{T}},\zeta\zeta^{\mathrm{T}}\rangle],\quad\forall i,j\in[n]\Big{\}}$
$=$ $\text{Pr}\Big{\{}\langle M_{i,j}M_{i,j}^{\mathrm{T}},\zeta\zeta^{\mathrm{T}}\rangle\cdot\|\zeta\|_{2}^{-2}\geq\varepsilon\cdot\mathbb{E}[\langle M_{i,j}M_{i,j}^{\mathrm{T}},\zeta\zeta^{\mathrm{T}}\rangle],\quad\forall i,j\in[n]\Big{\}}$
$\geq$ $\text{Pr}\Big{\{}\langle M_{i,j}M_{i,j}^{\mathrm{T}},\zeta\zeta^{\mathrm{T}}\rangle\geq\frac{\varepsilon}{\mu}\cdot\mathbb{E}[\langle M_{i,j}M_{i,j}^{\mathrm{T}},\zeta\zeta^{\mathrm{T}}\rangle],\quad\forall i,j\in[n],\quad\|\zeta\|_{2}^{-2}\geq\mu\Big{\}}$
$\geq$ $1-\sum_{(i,j)\in[n]}\text{Pr}\Big{\{}\langle M_{i,j}M_{i,j}^{\mathrm{T}},\zeta\zeta^{\mathrm{T}}\rangle<\frac{\varepsilon}{\mu}\cdot\mathbb{E}[\langle M_{i,j}M_{i,j}^{\mathrm{T}},\zeta\zeta^{\mathrm{T}}\rangle]\Big{\}}-\text{Pr}\Big{\{}\|\zeta\|_{2}>\frac{1}{\sqrt{\mu}}\Big{\}}$
$\geq$ $1-n^{2}\cdot\sqrt{\frac{\varepsilon}{\mu}}-\mu.$
As long as we take $\mu=33/100$, $\varepsilon=4\cdot(33/100)^{3}/n^{4}$, it holds that
$\text{Pr}\Big{\{}\langle M_{i,j}M_{i,j}^{\mathrm{T}},\widetilde{S}\rangle\geq\varepsilon\cdot\big{(}\widehat{f}_{i}+\widehat{g}_{j}\big{)},\quad\forall i,j\in[n]\Big{\}}\geq 1/100.$
In other words, there exists $\xi$ such that
$(\varepsilon\cdot\widehat{f}_{i})+(\varepsilon\cdot\widehat{g}_{j})\leq\langle M_{i,j}M_{i,j}^{\mathrm{T}},\xi\xi^{\mathrm{T}}\rangle,\quad\forall i,j\in[n]^{2},\qquad\|\xi\|_{2}=1.$
This indicates that $(\xi\xi^{\mathrm{T}},\varepsilon\cdot\widehat{f},\varepsilon\cdot\widehat{g})$ is a feasible solution to (KMS), and consequently,
[…]
$\square$
16 Rank Reduction Algorithm
In this section, we develop a rank reduction algorithm that, based on the near-optimal solution (denoted as $\widehat{S}$) returned from Algorithm 1, finds an alternative solution of the same (or smaller) optimality gap while satisfying the desired rank bound in Theorem 4.5.
Algorithm 4 Rank reduction algorithm for (SDR)
1:  Run Algorithm 1 to obtain $\delta$-optimal solution to (SDR), denoted as $\widehat{S}$.
// Step 2: Find $n$ binding constraints
2:  Run Hungarian algorithm [41] to solve OT (11) with $S\equiv\widehat{S}$, and obtain an optimal assignment $\sigma:~{}[n]\to[n]$ together with dual optimal solutoin $(f^{*},g^{*})$.
// Step 3-9: Calibrate low-rank solution using a greedy algorithm
3:  Initialize $\delta^{*}=1$
4:  while $\delta^{*}>0$ do
5:     Perform eigendecomposition $\widehat{S}=Q\Lambda Q^{\mathrm{T}}$, where $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{r})$ with $\text{rank}(\widehat{S})=r$
6:     Find a direction $Y=Q\Delta Q^{\mathrm{T}}$, where $\Delta\in\mathcal{S}^{r}$ is some nonzero matrix satisfying
$\text{Trace}(\Delta)=0,\langle Q^{\mathrm{T}}M_{i,\sigma(i)}M_{i,j}^{\mathrm{T}}Q,\Delta\rangle=0,\quad\forall i\in[n].$
7:     If such $Y$ does not exist, then break the loop.
8:     Take new solution $\widehat{S}(\delta^{*}):=\widehat{S}+\delta^{*}Y$, where
$\delta^{*}=\underset{\delta\geq 0}{\arg\max}~{}\Big{\{}\delta:~{}\lambda_{\min}(\Lambda+\delta\Delta)\geq 0\Big{\}}.$
9:     Update $\widehat{S}=\widehat{S}(\delta^{*})$
10:  end while
11:  Return $\widehat{S}$
Step (i): Find $n$ binding constraints. First, we fix $S\equiv\widehat{S}$ in (14) and finds the optimal solution $(f^{*},g^{*})$ such that only $n$ constraints out of $n^{2}$ constraints are binding. It suffices to apply the Huangarian algorithm [41] to solve the following balanced discrete OT problem
$\max_{f,g\in\mathbb{R}^{n}}\left\{\frac{1}{n}\sum_{i=1}^{n}(f_{i}+g_{i}):~{}~{}f_{i}+g_{j}\leq c_{i,j}\right\}=\min_{\pi\in\Gamma_{n}}~{}\left\{\sum_{i,j=1}^{n}\pi_{i,j}c_{i,j}\right\}$
where $c_{i,j}=\langle M_{i,j}M_{i,j}^{\mathrm{T}},\widehat{S}\rangle$. The output of the Huangarian algorithm is a deterministic optimal transport that moves $n$ probability mass points from the left marginal distribution of $\pi$ to the right, which is denoted as a bijection $\sigma$ that permutes $[n]$ to $[n]$. Thus, these $n$ binding constraints are denoted as
$f_{i}^{*}+g_{\sigma(i)}^{*}\leq\langle M_{i,j}M_{i,j}^{\mathrm{T}},S\rangle,\quad i\in[n].$
We denote by the intersection of these $n$ constraints as $\mathcal{E}_{n}$.
Step (ii): Calibrate low-rank solution using a greedy algorithm. Second, let us assume $\widehat{S}$ is not an extreme point of $\mathcal{S}_{2n}\cap\mathcal{E}_{n}$, since otherwise one can terminate the algorithm to output $\widehat{S}$ following the proof of Theorem 4.5. We run the following greedy rank reduction procedure:
  • (I)
    We find a direction $Y\neq 0$, along which $\widehat{S}$ remains to be feasible, and the null space of $\widehat{S}$ is non-decreasing.
  • (II)
    Then, we move $\widehat{S}$ along the direction $Y$ until its smallest non-zero eigenvalue vanishes. We update $\widehat{S}$ to be such a new boundary point.
  • (III)
    We terminate the iteration until no movement is available.
To achieve (I), denote the eigendecomposition of $\widehat{S}$ with $\text{rank}(\widehat{S})=r$ as
$\widehat{S}=\begin{pmatrix}Q&0\end{pmatrix}\begin{pmatrix}\Lambda&0\\ 0&0\end{pmatrix}\begin{pmatrix}Q^{\mathrm{T}}&0\end{pmatrix}=Q\Lambda Q^{\mathrm{T}}$
where $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{r})$ with $\lambda_{1}\geq\cdots\geq\lambda_{r}>0$ and $Q\in\mathbb{R}^{2n\times r}$. To ensure $\widehat{S}+\delta Y\in\mathcal{S}_{2n}\cap\mathcal{E}_{n}$ while $\text{Null}(\widehat{S}+\delta Y)\supseteq\text{Null}(\widehat{S})$, for some stepsize $\delta>0$, it suffices to take
$Y=\begin{pmatrix}Q&0\end{pmatrix}\begin{pmatrix}\Delta&0\\ 0&0\end{pmatrix}\begin{pmatrix}Q^{\mathrm{T}}&0\end{pmatrix}=Q\Delta Q^{\mathrm{T}},$
where $\Delta\in\mathcal{S}^{r}\setminus\{0\}$ is a symmetric matrix satisfying
$\text{Trace}(\Delta)=0,\quad\langle M_{i,j}M_{i,j}^{\mathrm{T}},Q\Delta Q^{\mathrm{T}}\rangle=0,\quad i\in[n].$
To achieve (II), it suffices to solve the one-dimensional optimization
$\delta^{*}=\underset{\delta\geq 0}{\arg\max}~{}\Big{\{}\delta:~{}\lambda_{\min}(\Lambda+\delta\Delta)\geq 0\Big{\}},$ (26)
where $\lambda_{\min}(\cdot)$ denotes the smallest eigenvalue of a given matrix. the optimization above admits closed-form solution update. Let eigenvalues of $\Delta$ be $\lambda_{1}^{\prime}\geq\cdots\geq\lambda_{r}^{\prime}$. It suffices to solve
$\delta^{*}=\underset{\delta\geq 0}{\arg\max}~{}\Big{\{}\delta:~{}\min_{i\in[r]}~{}(\lambda_{i}+\delta\lambda_{i}^{\prime})\geq 0\Big{\}}.$
As long as $\lambda_{r}^{\prime}\geq 0$, we return $\delta^{*}=0$. Otherwise, let $i$ be an index such that $\lambda_{i}^{\prime}\geq 0>\lambda_{i+1}^{\prime}$. We take $\delta^{*}=\max_{i<j\leq r}-\frac{\lambda_{j}}{\lambda_{j}^{\prime}}$ as the desired optimal solution.
The overall algorithm is summarized in Algorithm 4. Its performance guarantee is summarized in Propositions 17.1, 17.3, and Theorem 4.7.
17 Proof of Theorem 4.7
The proof of this theorem is separated into two parts.
Proposition 17.1
The rank of iteration points in Algorithm 4 strictly decreases. Thus, Algorithm 4 is guaranteed to terminate within $2n$ iterations.
Proof 17.2
Proof of Proposition 17.1. Assume on the contrary that $\text{rank}(\widehat{S}(\delta^{*}))=\text{rank}(\widehat{S})=r$. Since $\widehat{S}(\delta^{*})=Q(\Lambda+\delta^{*}\Delta)Q^{\mathrm{T}}$, the positive eigenvalues of $\widehat{S}(\delta^{*})$ are those of the matrix $\Lambda+\delta^{*}\Delta$. According to the solution structure of (26), this could happen only when $\Lambda+\delta^{*}\Delta\succ 0$, i.e., either $\delta^{*}=0$ or $\Delta\succeq 0$. For the first case, this algorithm terminates. For the second case, since $\text{Trace}(\Delta)=0,\Delta\in\mathcal{S}^{r}$, it implies that $\Delta=0$, which is a contradiction.
Thus, the rank of the iteration point reduces by at least $1$ in each iteration. $\square$
Proposition 17.3
Let $S^{*}$ be the output of Algorithm 4. Then, it holds that
  1. (I)
    $S^{*}$ is a $\delta$-optimal solution to (SDR).
  2. (II)
    The rank of $S^{*}$ satisfies
    $\text{rank}(S^{*})\leq 1+\left\lfloor\sqrt{2n+\frac{9}{4}}-\frac{3}{2}\right\rfloor.$
Proof 17.4
Proof of Proposition 17.3. Recall the solution $\widehat{S}$ obtained from Step 1 of Algorithm 4 satisfies
$F(\widehat{S})$ $=\min_{\pi\in\Gamma_{n}}~{}\sum_{i,j}\pi_{i,j}\langle M_{i,j}M_{i,j}^{\mathrm{T}},\widehat{S}\rangle$
[…]
Since Step 2 of Algorithm 4 solves the OT problem exactly, we obtain
[…]
Since Step 3-7 always finds feasible solutions to the $n$ binding constraints
$f_{i}^{*}+g_{\sigma(i)}^{*}\leq\langle M_{i,j}M_{i,j}^{\mathrm{T}},S\rangle,\quad i\in[n],$
for any iteration points from Step 3-7, denoted as $\widetilde{S}$, the pair $(\widetilde{S},f^{*},g^{*})$ is guaranteed to be the $\delta$-optimal solution to (14), a reformulation of (SDR). Hence we finish the proof of Part (I).
For the second part, assume on the contrary that $r=\text{rank}(S^{*})\geq 1+\left\lfloor\sqrt{2n+\frac{9}{4}}-\frac{3}{2}\right\rfloor$. It implies $n+1<r(r+1)/2$. Recall that Step 6 of Algorithm 4 essentially solves a linear system with $n+1$ constraints and $r(r+1)/2$ variables, so a nonzero matrix $\Delta$ is guaranteed to exist. Thus, one can pick a sufficiently small $\delta>0$ such that $\lambda_{\min}(\Lambda+\delta\Delta)\geq 0$, which contradicts to the termination condition $\delta^{*}=0$. Thus, we finish the proof of Part (II). $\square$
Combining both parts, we start to prove Theorem 4.7.
Proof 17.5
Proof of Theorem 4.7. Algorithm 4 satisfies the requirement of Theorem 4.7. For computational complexity, the computational cost of Step 2 of Algorithm 4 is $\mathcal{O}(n^{3})$. In each iteration from Step 3-7, the most computationally expansive part is to solve Step 6, which essentially solves a linear system with $n+1$ constraints and $r(r+1)/2$ variables. The conservative bound $r\leq 2n$. Hence, the worst-case computational cost of Step 6 (which can be achieved using Gaussian elimination) is
$\mathcal{O}((n+1+r(r+1)/2)\cdot(n+1)^{2})=\mathcal{O}(n^{4}).$
Since Algorithm 4 terminates within at most $2n$ iterations, the overall complexity of it is $\mathcal{O}(n^{5})$. $\square$
18 Numerical Implementation Details
18.1 Setup for Computing KMS Wasserstein Distance
When implementing our mirror ascent algorithm, for small sample size $(n\leq 200)$, we use the exact algorithm adopted from https://pythonot.github.io/ to solve the inner OT; whereas for large sample size, we use the approximation algorithm adopted from https://github.com/YilingXie27/PDASGD to solve this subproblem. For the baseline BCD approach, we implement it using the code from github.com/WalterBabyRudin/KPW_Test/tree/main.
18.2 Setup for High-dimensional Hypothesis Testing
For baselines Sinkhorn Div, SW, MS, we implement them by calling the well-established package POT [20]. For the ME baseline, we implement it using the code from https://github.com/wittawatj/interpretable-test.
Next, we outline how to generate the datasets for two-sample testing experiments in Fig. 4:
  1. (I)
    The Gaussian covariance shift dataset was generated by taking
    $\mu=\mathcal{N}(0,I_{d}),\nu=\mathcal{N}(0,I_{d}+\rho E),$
    where the dimension $d=200$, the sample size $n=200$, and $E$ is the all-one matrix. We vary the hyper-parameter $\rho$ from 0 to $0.06$.
  2. (II)
    The Gaussian mixture dataset was generated by taking
    $\mu=\frac{1}{2}\mathcal{N}(0_{d},I_{d})+\frac{1}{2}\mathcal{N}(0.5\cdot 1_{d},I_{d}),\quad\nu=\frac{1}{2}\mathcal{N}(0_{d},I_{d}+0.05E)+\frac{1}{2}\mathcal{N}(0.5\cdot 1_{d},I_{d}+0.05E),$
    where the dimension $d=40$, and we vary the sample size $n$ from 100 to 700.
  3. (III)
    For MNIST or CIFAR10 examples, we take $\mu$ as the uniform distribution subsampled from the target dataset, denoted as $\mu=p_{\text{data}}$. We take $\nu$ as the one having a change of abundance, i.e., $\nu=0.85p_{\text{data}}+0.15p_{\text{data of label 1}}$, where $p_{\text{data of label 1}}$ corresponds to the distribution of a subset of the class with label $1$.
Finally, we outline the procedure for practically using the KMS Wasserstein distance for two-sample testing, based on the train-test split method:
  1. (I)
    We first do the $50\%$-$50\%$ training-testing data split such that $x^{n}=x^{\text{Tr}}\cup x^{\text{Te}}$ and $y^{n}=y^{\text{Tr}}\cup y^{\text{Te}}$.
  2. (II)
    Then we compute the optimal nonlinear projector $f$ for training data $(x^{\text{Tr}},y^{\text{Tr}})$. We specify the testing statistic as the Wasserstein distance between projected testing data $(f_{\#}x^{\text{Te}},f_{\#}y^{\text{Te}})$.
  3. (III)
    Then we do the permutation bootstrap strategy that shuffles $(x^{\text{Te}},y^{\text{Te}})$ for many times, e.g., $L=500$ times. For each time $t$, the permuted samples are $(x^{\text{Te}}_{(t)},y^{\text{Te}}_{(t)})$, and we obtain the testing statistic as the Wasserstein distance between projected samples (using the estimated projector $f$, denoted as $\mathcal{W}({f}_{\#}x^{\text{Te}}_{(t)},{f}_{\#}y^{\text{Te}}_{(t)})$.
  4. (IV)
    Finally, we obtain the threshold as $(1-\alpha)$-quantile of the testing statistics for permuted samples, where the type-I error $\alpha=0.05$.
18.3 Setup for Human Activity Detection
The MSRC-12 Kinect gesture dataset contains sequences of human body movements recorded by 20 sensors collected from 30 users performing 12 different gestures. We pre-process the dataset by extracting 10 different users such that in the first 600 timeframes, they are performing the throwing action, and in the remaining 300 timeframes, they are performing the lifting action.
18.4 Setup for Generative Modeling
We follow Deshpande et al. [16] to design the optimization algorithm that solves the problem $\min_{\theta}~{}\mathcal{D}\big{(}p_{\text{data}},(f_{\theta})_{\#}p_{\text{noise}}\big{)},$ where $p_{\text{data}}$ denotes the empirical distribution of MNIST dataset, $p_{\text{noise}}$ denotes the Gaussian noise, and $(f_{\theta})_{\#}p_{\text{noise}}$ represents the distribution of the fake image dataset. We specify $f_{\theta}$ as a $4$-layer feed-forward neural-net with leaky relu activation, and $\theta$ denotes its weight parameters. We train the optimization algorithm in $30$ epoches. From Fig. 6, we observe that the KMS Wasserstein distance provides fake images that are more closer to the ground truth compared with the Sliced Wasserstein distance.
19 Additional Numerical Study
19.1 Experiment on Significance Level
Here we add the experiment in the following table to show that as long as we take $\alpha=0.05$, the practical type-I error of KMS Wasserstein test is guaranteed to be controlled within 0.05.
Table 3: Type-I Error for two-sample testing with Gaussian mixture dataset.
Method 20 40 80 160 180 200
KMS $0.061\pm 0.015$ $0.055\pm 0.012$ $0.043\pm 0.014$ $0.059\pm 0.011$ $0.042\pm 0.014$ $0.062\pm 0.015$
19.2 Rank Reduction Algorithm
Recall that Theorem 4.5 provides the rank bound regarding some optimal solution from SDR. In this part, we compare the rank of the matrix estimated from Algorithm 1 with our theoretical rank bound based on the CIFAR10 dataset. For a given positive semidefinite matrix, we calculate the rank as the number of eigenvalues greater than the tolerance 1e-6. The numerical performance is summarized in Table 4. In these cases, one can run our rank reduction algorithm to obtain a low-rank solution. Fig. 8 illustrates the procedure by showing the probability mass values associated with the eigenvectors of the estiamted solution $\widehat{S}$. From the plot, we find that our rank reduction algorithm is capable of producing low-rank solutions even though the matrix size is large.
Figure 8: Procedure of rank reduction algorithm for CIFAR10 example with sample size $n=200$.
Table 4: Numerical performance on rank for CIFAR10 dataset
Sample Size $n$ Rank Obtained from Algorithm 1 Rank Bound from Theorem 4.5
200 400 19
250 500 21
300 600 24
350 700 26
400 800 27
450 900 29
500 1000 31